### v1

# Existence of Solutions to the Seiberg-Witten Vortex Equations with Exponential Decay on the Plane

Abstract.
Clifford Taubes showed that the moduli space of the variational equation of the Yang-Mills-Higgs functional on the plane is non-empty, and its elements correspond to "vortices". Inspired by this result, in this paper, we show that the moduli space of the Hitchin-type dimensional reduction of the Seiberg-Witten equations on the plane contains both exponentially decayed solutions and polynomial growth solutions. Furthermore, we show that there is correspondence from the moduli space of exponentially decayed and polynomial growth solutions to the symmetric products of complex numbers. The correspondence restricted to the latter is a surjective map.
Keywords: Vortex equations, Yang-Mills-Higgs functional, Seiberg-Witten equations, Vekua equations, Generalized analytic functions, Gauge theory
2020 Mathematics Subject Classification:
Primary 53Cxx, 57Rxx, 58Jxx, 57Kxx, 30G20, 30Cxx
1. Introduction
1.1. Main Results
In this paper, we study some analytic aspects of the solutions of a dimensional reduction to the Seiberg-Witten equations on $\mathbf{R}^{2}$. The Seiberg-Witten equations were first introduced by Seiberg and Witten in [seiberg1994electric, seiberg1994monopoles]. These equations come from gauge theory of Mathematical Physics. Even though they are physically motivated, the precise mathematical interpretation of gauge fields and matter fields as connections and sections on a vector bundle in differential geometry has turned the equations into one of the most instrumental tools in studying problems in low-dimensional topology and symplectic topology (see, e.g., [MR1339810, taubes2000seiberg, kronheimer2007monopoles, MR1367507] for some survey).
The dimensional reduction of the Seiberg-Witten equations we consider here is the one where solutions are assumed to be translational invariant in the last two coordinates of $\mathbf{R}^{4}$. Effectively, the newly derived equations (which we will refer to as the Seiberg-Witten vortex equations) consist of objects defined only in two dimensions (cf. (3.12), (4.1)), where they are known to be conformally invariant. Thus, they can also be considered globally on a Riemann surface. Nevertheless, in this paper, we choose not to discuss the global aspect of the equations and focus only on the local analysis. The global aspect of the equations was discussed in Dey’s work [MR1952133, MR2698453], where it is shown that the moduli space of the Seiberg-Witten vortex equations is a symplectic, almost complex manifold that also has a hyperKähler structure.
The derivation of the Seiberg-Witten vortex equations is inspired by the work of Hitchin. In [MR0887284], Hitchin studied a class of solutions to the (anti)self-dual Yang-Mills equations on $\mathbf{R}^{4}$ that are translational invariant in the last two coordinates. The resulting equations on $\mathbf{R}^{2}$ are now called the Hitchin equations. The Hitchin moduli space on a Riemann surface has many striking properties, one of which is the correspondence with the moduli space of Higgs bundle [hitchin1987stable]. The exploitation of this correspondence led to many applications in number theory, notably the work of Ngô in the Langlands program [ngo2006fibration]. For this reason, it seems interesting and potentially fruitful to apply the same pathway to the Seiberg-Witten equations, as already initiated by Dey’s work [MR1952133, MR2698453]. Note that the dimensional reduction of this type is important in establishing a Topological Quantum Field Theory (TQFT) of dimension $(2+1)$.
We will refrain from making any more comments about the global geometric aspect of the various gauge theoretic equations mentioned above. From this point on, we will only focus on the analysis of the Seiberg-Witten vortex equations. Note that the Seiberg-Witten equations are known not to have interesting $L^{2}$ solutions on $\mathbf{R}^{4}$. However, on $\mathbf{R}^{2}$ (and Riemann surfaces), the Seiberg-Witten vortex equations do have interesting solutions. As a system of PDEs, in plain terms, the Seiberg-Witten vortex equations look for the unknowns $(A_{0},A_{1},\psi_{1},\psi_{2})$ that satisfy (cf. (4.1))
$\begin{cases}2\dfrac{\partial\psi_{2}}{\partial\overline{z}}+i(A_{0}+iA_{1})\psi_{2}=0,\\ 2\dfrac{\partial\psi_{1}}{\partial z}+i(A_{0}-iA_{1})\psi_{1}=0,\\ i\left(\dfrac{\partial A_{1}}{\partial x_{0}}-\dfrac{\partial A_{0}}{\partial x_{1}}\right)=\dfrac{i}{2}(|\psi_{1}|^{2}-|\psi_{2}|^{2}).\end{cases}$
Here $A_{0},A_{1}$ are real-valued functions on $\mathbf{R}^{2}$, and $\psi_{1},\psi_{2}$ are complex-valued functions on $\mathbf{R}^{2}$. The Seiberg-Witten vortex equations have a symmetry given by the gauge group $\mathcal{G}=Maps(\mathbf{R}^{2},U(1))$, i.e, the solutions are invariant under the action of the gauge group $\mathcal{G}$. By moduli space of the equations, we mean the space of solutions quotient out by the $\mathcal{G}$-symmetry. It is not difficult to write down a solution of (4.1), albeit "trivial". Here is one: $(\partial f/\partial x_{0},\partial f/\partial x_{1},0,0)$, where $f$ is any smooth function on $\mathbf{R}^{2}$. In fact, one can be a little bit algebraically creative and realize that
Theorem 1.1 (cf. Proposition 3.9, Corollary 3.10).
For any $(c_{1},c_{2})\in\mathbf{C}^{*}\times\mathbf{C}^{*}$ and $\theta\in\mathbf{R}$, $(A_{0},A_{1},\psi_{1},\psi_{2})$ given by
$(-2c_{2},0,c_{1}e^{i\theta}(\overline{z}-\overline{z_{1}})^{n_{1}}\cdots(\overline{z}-\overline{z_{k}})^{n_{k}}e^{ic_{2}(z+\overline{z})},c_{1}(z-z_{1})^{n_{1}}\cdots(z-z_{k})^{n_{k}}e^{ic_{2}(z+\overline{z})})$
is always a solution of (4.1), where $\mathbf{C}^{*}$ denotes $\mathbf{C}\setminus\{0\}$.
Denote by $Vor_{\mathfrak{p}}(\mathbf{C})$ the moduli space of solutions to (4.1) of the type in Theorem 1.1. Consequently, we immediately have the following theorem.
Theorem 1.2 (cf. Theorem 3.11).
There is a surjective map $\eta_{\mathfrak{p}}:Vor_{\mathfrak{p}}(\mathbf{C})\to\bigcup_{n\in\mathbf{N}}Sym^{n}(\mathbf{C})$.
Note that the solutions given in Theorem 1.1 and Theorem 1.2 are of polynomial growth and the "connection" $A=iA_{0}dx_{0}+iA_{1}dx_{1}$ is always flat. Thus, it is natural to ask if there are any other types of solutions to (4.1) that exhibit different behaviors at infinity and the connection part is not necessarily flat, e.g. exponentially decayed solutions. Aside from the geometric meaning of the equations, the existence of such solutions is an interesting analysis problem in its own right. The main result of the paper shows exactly that.
Theorem 1.3 (cf. Theorem 4.1).
Let $C$ be any positive number. There exists a smooth solution $(A_{0},A_{1},\psi_{1},\psi_{2})$ of (4.1) with the property (E) that $\psi_{1},\psi_{2}$ have the same argument and that for every $\epsilon>0$, there exists a $c>0$ such that $C-|\psi_{2}|^{2}\leq c\cdot\exp(-(1-\epsilon)|x|)$.
The heart of the analysis that goes into the proof of Theorem 1.3 is the existence and uniqueness result of a certain Kazdan-Warner equation with an appropriate boundary condition.
Theorem 1.4 (cf. Theorem 4.4).
For each $C>0$, the equation $-\Delta u+4Ce^{u}-2C=-4\pi\displaystyle\sum_{k=1}^{n}\delta(z-z_{k})$ has a real analytic solution with the condition that $u\leq 0$ and $u\to 0$ in a certain appropriate sense. Here $\delta$ denotes the Dirac measure.
The exponential decay in Property (E) seems to be out of place. However, it is quite natural to consider. The Seiberg-Witten vortex equations can be thought of as a variant of the vortex equations which were introduced by Ginzburg and Landau to study the theory of superconductivity. From the Mathematics perspective, they are the absolute minimum condition for the Yang-Mills-Higgs functional (cf. Subsection 3.1). Many authors have studied the existence of solutions related to these types of vortex equations derived from Yang-Mills-Higgs models (see, e.g, [MR0614447, MR0573986, garcia1994direct, bradlow1990vortices, bradlow1991special, MR4516074]). Using the Yang-Mills-Higgs functional, the solutions of the Ginzburg-Landau vortex equations can be shown to have exponential decay a priori. To the best of our knowledge, we do not know whether our (4.1) is also the absolute minimum condition for some Yang-Mills-Higgs-type functional. Regardless, we note that by Theorem 1.1, Theorem 1.2 and Theorem 1.3, solutions of (4.1) can exhibit both polynomial growth or exponential decay, a feature that is not shared with the Ginzburg-Landau vortex equations.
As a final remark, the first two equations of the Seiberg-Witten vortex equations can be thought of as a system of Vekua equations. This insight is crucial for us to study the zeroes of exponential decay solutions of (4.1). We develop the basic theory of solutions to these systems of this type that we require and associate the functions that solve these systems with the solutions of vortex equations. This is the first time that a system of Vekua equations of this kind (to be precise a Vekua equation and the result of applying complex conjugation to both sides of a Vekua equation) has been studied. This connection between the classical complex analysis structure of Vekua equations and the gauge equations of mathematical physics is novel. See [Vek], [Bers], and Section 2 for background on Vekua equations.
1.2. Organization
The paper is organized as follows. In Section 2, we present some material about the classical Vekua equation and prove some new solution representation results about the system of Vekua-type equations that we use later. Section 3 contains background about gauge theory and introduces the terminology of the main results of the paper. In Section 4, we prove the main results using a technique from calculus of variations. In Section 5, we give some brief comments about the limitations of the technique we used when applied to the situation of the Seiberg-Witten vortex equations with Higgs fields (cf. (3.12)) and discuss some future problems of extending the technique to other generalizations of the Seiberg-Witten equations.
Acknowledgement.
Parts of this work were carried out during the first-named author’s visit to Washington University in St. Louis in the Fall of 2023. The authors would like to thank the generous accommodation of the university. The second-named author is also grateful for many conversations with Aliakbar Daemi regarding various vortex equations on Riemann surfaces. ∎
2. Preliminaries
2.1. Classical Vekua Equation
In this section, we provide some background about nonhomogeneous Cauchy-Riemann equations that will be used throughout. We work on the open unit disk $D$ of the complex plane. A nonhomogeneous Cauchy-Riemann equation is any equation of the form
(2.1) $\dfrac{\partial w}{\partial\overline{z}}=f,$
where $f\not\equiv 0$. To study solutions to equations of the form of (2.1), we first recall the classic Cauchy-Pompieu theorem.
Theorem 2.1 ($\bar{\partial}$-Poincare Lemma [MR0507725]; Cauchy-Pompeiu Theorem, Theorem 20 [BegBook] ).
Every $w\in C^{1}(D)\cap C(\overline{D})$ has the representation
[…]
where $\zeta=\eta+i\xi$.
Since $w\in C(\partial D)$ in the hypothesis of the last theorem, it follows that the contour integral on the right-hand side of (2.2) is a holomorphic function (see [GK]). By applying $\dfrac{\partial}{\partial\overline{z}}$ to both sides of (2.2), we have
[…]
Hence, the area integral is a right-inverse to the Cauchy-Riemann operator $\dfrac{\partial}{\partial\overline{z}}$. This behavior persists for rougher classes of functions than those considered in Theorem 2.1, as the next theorem shows.
Theorem 2.2 (Theorem 1.16 [Vek]).
For $f\in L^{1}(D)$, every solution of
$\dfrac{\partial w}{\partial\overline{z}}=f$
has the form
$w=\varphi+T(f),$
where
[…]
and $\zeta=\eta+i\xi$.
A well known property of the operator $T(\cdot)$ defined in the last theorem is the following.
Theorem 2.3 (Theorem 1.19 [Vek]).
For every $f\in L^{q}(D)$, $q>2$, $T(f)\in C^{0,\alpha}(\overline{D})$, where $\alpha=\dfrac{q-2}{q}$.
A well-studied nonhomogeneous Cauchy-Riemann equation is the Vekua equation
(2.3) $\dfrac{\partial w}{\partial\overline{z}}=Aw+B\overline{w},$
where $A,B$ are functions in a Lebesgue space on $D$. Solutions of this equation were classically studied by I. N. Vekua [Vek] (who called them generalized analytic functions) in their study of infinitesimal bendings of surfaces and L. Bers [Bers] (who called them pseudoanalytic functions) in their study of functions that generalize holomorphic functions. The Vekua equation (2.3) is an important class of nonhomogeneous Cauchy-Riemann equations because its solutions share many properties of holomorphic functions. This similarity is realized by the following representation formula.
Theorem 2.4 (“The Basic Lemma” [Vek]).
Every function $w$ that solves
$\dfrac{\partial w}{\partial\overline{z}}=Aw+B\overline{w}$
in $D$, where $A,B\in L^{q}(D)$, $q>2$, has the form
$w=\varphi e^{\phi},$
where $\varphi$ is holomorphic in $D$ and
$\phi(z):=\begin{cases}T\left(A+B\dfrac{\overline{w}}{w}\right)(z),&w(z)\neq 0\\ 0,&w(z)=0\end{cases}.$
This representation is called the “similarity principle” or the “representation of the first kind.” From this representation, we see that generalized analytic functions inherit their zero set behavior from holomorphic functions, and since $\phi\in C^{0,\alpha}(\overline{D})$ by Theorem 2.3, it follows that $|e^{\phi}|$ is bounded above and below away from zero, so many other results about holomorphic functions that rely solely on size estimates are recoverable for generalized analytic functions.
Equations of the form
$\dfrac{\partial w}{\partial\overline{z}}=Aw+B\overline{w}+f,$
where $f\not\equiv 0$, are called nonhomogeneous Vekua equations, and by Theorem 2.2, if $Aw+B\overline{w}+f\in L^{1}(D)$, then
$w=\varphi+T(Aw+B\overline{w}+f),$
for some holomorphic function $\varphi$. However, we lose the similarity principle representation. In general, there is no reason that $\dfrac{f}{w}\in L^{q}(D)$, $q>2$, so $T\left(A+B\dfrac{\overline{w}}{w}+\dfrac{f}{w}\right)$ may not converge. In the special case that $\dfrac{f}{w}\in L^{q}(D)$, $q>2$, we have the following.
Proposition 2.5.
For $A,B\in L^{q}(D)$, $q>2$, any solution $w$ of
$\dfrac{\partial w}{\partial\overline{z}}=Aw+B\overline{w}+f,$
such that $\dfrac{f}{w}\in L^{q}(D)$, $q>2$, has the representation
$w=\varphi e^{\phi},$
where $\varphi$ is holomorphic and $\phi$ defined as
$\phi(z):=\begin{cases}T\left(A+B\dfrac{\overline{w}}{w}+\dfrac{f}{w}\right)(z),&w(z)\neq 0\\ 0,&w(z)=0\end{cases}$
is in $C^{0,\alpha}(\overline{D})$, $\alpha=\dfrac{q-2}{q}$.
Proof.
Let $w$ be a solution of
$\dfrac{\partial w}{\partial\overline{z}}=Aw+B\overline{w}+f,$
such that $\dfrac{f}{w}\in L^{q}(D)$, $q>2$. Since $T(\cdot)$ is a right-inverse to $\dfrac{\partial}{\partial\overline{z}}$ by Theorem 2.2, it follows that
$\dfrac{\partial}{\partial\overline{z}}\left(\dfrac{w}{e^{\phi}}\right)$ $=\dfrac{e^{\phi}\dfrac{\partial w}{\partial\overline{z}}-w\dfrac{\partial}{\partial\overline{z}}(e^{\phi})}{(e^{\phi})^{2}}$
$=\dfrac{e^{\phi}(Aw+B\overline{w}+f)-we^{\phi}\left(A+B\dfrac{\overline{w}}{w}+\dfrac{f}{w}\right)}{(e^{\phi})^{2}}$
$=0.$
Hence, $\dfrac{w}{e^{\phi}}$ is holomorphic. Since $A+B\dfrac{\overline{w}}{w}+\dfrac{f}{w}\in L^{q}(D)$, $q>2$, it follows that $\phi\in C^{0,\alpha}(D)$, by Theorem 2.3
The hypothesis of the last proposition is very strict. In general, this result would be of little interest. However, in the context of vortex equations, there are situations where this hypothesis is natural, and it allows us to extend the most useful of representations of functions of this type to the nonhomogeneous Vekua equations.
2.2. System of Vekua Equations
Next, we consider systems of the form
(2.4) $\begin{cases}\dfrac{\partial w_{1}}{\partial z}+iAw_{1}+Bw_{2}=0\\ &\\ \dfrac{\partial w_{2}}{\partial\overline{z}}+i\overline{A}w_{2}+\overline{B}w_{1}=0\end{cases},$
where $A$ and $B$ are functions.
Note that system (2.4) is comprised of a nonhomogeneous Vekua equation and the result of applying complex conjugation to both sides of a nonhomogeneous Vekua equation. We work to analyze pairs of solutions $(w_{1},w_{2})$ to systems in the form of (2.4).
Theorem 2.6.
Let $A\in L^{q}(D)$, $q>2$. Every solution pair $(w_{1},w_{2})$ of the system
$\begin{cases}\dfrac{\partial w_{1}}{\partial z}+iAw_{1}+Bw_{2}=0\\ &\\ \dfrac{\partial w_{2}}{\partial\overline{z}}+i\overline{A}w_{2}+\overline{B}w_{1}=0,\end{cases}$
such that $B\dfrac{w_{2}}{w_{1}},\overline{B}\dfrac{w_{1}}{w_{2}}\in L^{q}(D)$, $q>2$, has the representation
$(w_{1},w_{2})=(\,\overline{e^{\phi_{1}}\varphi_{1}},e^{\phi_{2}}\varphi_{2})$
where
$\phi_{1}(z)=\begin{cases}T\left(-\left(\overline{iA+B\dfrac{w_{2}}{w_{1}}}\right)\right)(z),&w_{1}(z)\neq 0\\ 0,&w_{1}(z)=0\end{cases}$
and
$\phi_{2}(z)=\begin{cases}T\left(-\left(i\overline{A}+\overline{B}\dfrac{w_{1}}{w_{2}}\right)\right)(z),&w_{2}(z)\neq 0\\ 0,&w_{2}(z)=0\end{cases}$
are in $C^{0,\alpha}(\overline{D})$ and $\varphi_{1},\varphi_{2}$ are holomorphic.
Proof.
The system
$\begin{cases}\dfrac{\partial w_{1}}{\partial z}+iAw_{1}+Bw_{2}=0\\ &\\ \dfrac{\partial w_{2}}{\partial\overline{z}}+i\overline{A}w_{2}+\overline{B}w_{1}=0,\end{cases}$
is equivalent to
$\begin{cases}\dfrac{\partial w_{1}}{\partial z}=-(iAw_{1}+Bw_{2})\\ &\\ \dfrac{\partial w_{2}}{\partial\overline{z}}=-(i\overline{A}w_{2}+\overline{B}w_{1})\end{cases}.$
We now consider each of the equations individually. Observe that if
$\dfrac{\partial w_{1}}{\partial z}=-(iAw_{1}+Bw_{2}),$
then
$\dfrac{\partial\overline{w_{1}}}{\partial\overline{z}}=-(\,\overline{iAw_{1}+Bw_{2}}).$
Since $\left|\overline{B}\dfrac{\overline{w_{2}}}{\overline{w_{1}}}\right|=\left|B\dfrac{w_{2}}{w_{1}}\right|$ and $B\dfrac{w_{2}}{w_{1}}\in L^{q}(D)$, $q>2$, it follows that $\overline{B}\dfrac{\overline{w_{2}}}{\overline{w_{1}}}\in L^{q}(D)$, $q>2$. Hence, by Proposition 2.5,
$\overline{w_{1}}=e^{\phi_{1}}\varphi_{1},$
where
$\varphi_{1}(z):=\begin{cases}T\left(-\left(\overline{iA+B\dfrac{w_{2}}{w_{1}}}\right)\right)(z),&w_{1}(z)\neq 0\\ 0,&w(z)=0\end{cases}$
and $\varphi_{1}$ is holomorphic. Thus,
$w_{1}=\overline{e^{\phi_{1}}\varphi_{1}}.$
Similarly, since $\overline{B}\dfrac{w_{1}}{w_{2}}\in L^{q}(D)$, $q>2$, by hypothesis, it follows by Proposition 2.5 that
$w_{2}=e^{\phi_{2}}\varphi_{2},$
where
$\varphi_{2}(z)=\begin{cases}T\left(-\left(i\overline{A}+\overline{B}\dfrac{w_{1}}{w_{2}}\right)\right)(z),&w_{2}(z)\neq 0\\ 0,&w_{2}(z)=0\end{cases}$
and $\varphi_{2}$ is holomorphic. ∎
3. Vortex Equations
This section presents a brief introduction to mathematical gauge theory and sets up relevant terminologies that go into the statement of the main result of the paper. For more details, we direct the readers to the following references [MR0614447, MR0573986, MR3012377, MR1367507, MR1339810, MR1306021, MR2698453, MR0887284, SACLIOGLU1997675, MR1362874].
3.1. Gauge Theory Set-Up
From the physical perspective of Yang-Mills-Higgs theory in the $n$-dimensional Euclidean space $\mathbf{R}^{n}$, the variables are given by
  • A gauge potential $A=\sum_{j=1}^{n}A_{j}(x)dx_{j}$. In differential geometry, gauge potentials are to be understood as connections of certain principal $G$-bundle $P$ over $\mathbf{R}^{n}$, where $G$ is a Lie group. Thus, $A_{j}$ are functions defined on $\mathbf{R}^{n}$ that take values in the Lie algebra $\mathfrak{g}$ of $G$.
  • A matter field $\phi=\phi(x)$. Once again, from the mathematical perspective, $\phi$ should be thought of as a section of an associated vector bundle $E=P\times_{G}V\to\mathbf{R}^{n}$ that is defined as long as one has a representation $\rho:G\to GL(V)$, where $V$ is some finite dimensional vector space.
Here, $x$ denotes a point in $\mathbf{R}^{n}$, written in terms of the standard coordinate system as $x=(x_{1},\cdots,x_{n})$. Let $\{e_{1},\ldots,e_{n}\}$ be the standard orthonormal basis of $\mathbf{R}^{n}$. $G$ is called the gauge group, i.e, it is the group of transformations of the internal symmetry space $E$ that $\phi$ takes value in. The interaction between a gauge potential and matter field is via the notion of taking a covariant derivative. In particular,
$\nabla_{A}\phi=\displaystyle\sum_{j=1}^{n}(d\phi(e_{j})+\rho(A_{j})(\phi))dx_{j}.$
One should think of $\nabla_{A}\phi$ as an $E$-valued $1$-form defined on $\mathbf{R}^{n}$. A connection $A$ also determines for us the notion of curvature $F(A)$, which locally can be written as
$F(A)=dA+A\wedge A=\displaystyle\sum_{j,k}\dfrac{1}{2}\left(\dfrac{\partial A_{k}}{\partial x_{j}}-\dfrac{\partial A_{j}}{\partial x_{k}}+[A_{j},A_{k}]\right)dx_{j}\wedge dx_{k}.$
In what follows, we consider the case where $n=2$, the gauge group $G=U(1)$, $P$ is taken to be the trivial principal bundle $P=\mathbf{R}^{2}\times U(1)$, and $\rho:U(1)\to GL(\mathbf{C})$ to be the standard representation of $U(1)$ on $\mathbf{C}$ given by the (complex) scalar multiplication. As a result, the associated vector bundle $E$ simplifies to be the trivial complex line bundle $\mathbb{L}$. Since we are working on Euclidean space $\mathbf{R}^{2}$ where there is no interesting topology and the Lie algebra of $U(1)$ is simply $i\mathbf{R}$, various actors defined above can be simplified as follows. The matter field $\phi$ is now simply a $\mathbf{C}$-valued function on $\mathbf{R}^{2}$. The gauge potential $A$ is a purely imaginary-valued $1$-form on $\mathbf{R}^{2}$, written as $A=iA_{1}dx_{1}+iA_{2}dx_{2}$. Hence, the curvature $F(A)$ would just be the curl of $A$ given by
$F(A)=i\left(\dfrac{\partial A_{2}}{\partial x_{1}}-\dfrac{\partial A_{1}}{\partial x_{2}}\right)dx_{1}\wedge dx_{2}.$
Whereas, $\nabla_{A}\phi$ is now a complex valued $1$-form on $\mathbf{R}^{2}$.
Denote by $\mathcal{C}=i\Omega^{1}(\mathbf{R}^{2})\times C^{\infty}(\mathbf{R}^{2},\mathbf{C})$ the configuration space. Let $dVol$ be the standard Lebesgue measure on $\mathbf{R}^{2}$. The Euclidean Yang-Mills-Higgs action functional on $\mathcal{C}$ is given by
[…]
The various point-wise norms that appear in the integrand above deserve some justification. Firstly, since $\phi$ is complex-valued, $|\phi|^{2}=\phi\overline{\phi}$. Next, since $\nabla_{A}\phi=\sum_{j=1}^{2}(d\phi(e_{j})+iA_{j}\phi)dx_{j}$ is a $\mathbf{C}$-valued $1$-form on $\mathbf{R}^{2}$, we simply define $|\nabla_{A}\phi|^{2}$ to be the sum of the squares of the norm of the $\mathbf{C}$-component of the form. Similarly,
$|F(A)|^{2}=\left|\dfrac{\partial A_{2}}{\partial x_{1}}-\dfrac{\partial A_{1}}{\partial x_{2}}\right|^{2}.$
For convenience, from now on, we shall write $\partial_{j}:=\partial/\partial x_{j}$.
Lemma 3.1 (Bogomolny).
Let $\phi$ be written as $\phi_{1}+i\phi_{2}$ when we view it as a complex-valued function defined on $\mathbf{R}^{2}$ or $\begin{bmatrix}\phi_{1}&\phi_{2}\end{bmatrix}^{T}$ when viewed as a map from $\mathbf{R}^{2}\to\mathbf{R}^{2}$, where $\phi_{j}$ are $\mathbf{R}$-valued functions. We can re-write the Yang-Mills-Higgs action functional (3.1) as follows
$\mathscr{Y}(A,\phi)=\dfrac{1}{2}\iint_{\mathbf{R}^{2}}$ $[(\partial_{1}\phi_{1}-A_{1}\phi_{2})+(\partial_{2}\phi_{2}+A_{2}\phi_{1})]^{2}+$
$+[(\partial_{1}\phi_{2}+A_{1}\phi_{1})-(\partial_{2}\phi_{1}-A_{2}\phi_{2})]^{2}+$
$+\dfrac{1}{2}\iint_{\mathbf{R}^{2}}((\partial_{1}A_{2}-\partial_{2}A_{1})/2+|\phi|^{2}-1)^{2}+\dfrac{1}{2}\iint_{\mathbf{R}^{2}}F(A)+$
$+\iint_{\mathbf{R}^{2}}d(\phi_{2}(d\phi_{1}-iA\phi_{2})-\phi_{1}(d\phi_{2}+iA\phi_{1})).$
Proof.
When we view $\phi=\begin{bmatrix}\phi_{1}&\phi_{2}\end{bmatrix}^{T}:\mathbf{R}^{2}\to\mathbf{R}^{2}$, its total derivative in matrix form is written as
$d\phi=\begin{bmatrix}\partial_{1}\phi_{1}&\partial_{2}\phi_{1}\\ \partial_{1}\phi_{2}&\partial_{2}\phi_{2}\end{bmatrix}.$
As a result, $d\phi(e_{1})=\partial_{1}\phi_{1}+i\partial_{1}\phi_{2}$ and $d\phi(e_{2})=\partial_{2}\phi_{1}+i\partial_{2}\phi_{2}$. Thus, from the definition of the Yang-Mills-Higgs functional, we can re-write
$\mathscr{Y}(A,\phi)=\dfrac{1}{2}\iint_{\mathbf{R}^{2}}$ $\dfrac{1}{4}(\partial_{1}A_{2}-\partial_{2}A_{1})^{2}+|(\partial_{1}\phi_{1}-A_{1}\phi_{2})+i(\partial_{1}\phi_{2}+A_{1}\phi_{1})|^{2}+$
$+|(\partial_{2}\phi_{1}-A_{2}\phi_{2})+i(\partial_{2}\phi_{2}+A_{2}\phi_{1})|^{2}+(|\phi|^{2}-1)^{2}.$
By completion of squares, we can re-arrange the above as
$\mathscr{Y}(A,\phi)=$ $\dfrac{1}{2}\iint_{\mathbf{R}^{2}}[(\partial_{1}\phi_{1}-A_{1}\phi_{2})+(\partial_{2}\phi_{2}+A_{2}\phi_{1})]^{2}+$
$+\dfrac{1}{2}\iint_{\mathbf{R}^{2}}[(\partial_{1}\phi_{2}+A_{1}\phi_{1})-(\partial_{2}\phi_{1}-A_{2}\phi_{2})]^{2}+$
$+\dfrac{1}{2}\iint_{\mathbf{R}^{2}}((\partial_{1}A_{2}-\partial_{2}A_{1})/2+|\phi|^{2}-1)^{2}+\dfrac{1}{2}\iint_{\mathbf{R}^{2}}F(A)+$
$+\dfrac{1}{2}\iint_{\mathbf{R}^{2}}-(\partial_{1}A_{2}-\partial_{2}A_{1})|\phi|^{2}+$
$+\dfrac{1}{2}\iint_{\mathbf{R}^{2}}-2(\partial_{1}\phi_{1}-A_{1}\phi_{2})(\partial_{2}\phi_{2}+A_{2}\phi_{1})+$
$+\dfrac{1}{2}\iint_{\mathbf{R}^{2}}2(\partial_{1}\phi_{2}+A_{1}\phi_{1})(\partial_{2}\phi_{1}-A_{2}\phi_{2}).$
We simplify the integrand of the last three terms of the equation above as follows
$-(\partial_{1}A_{2}-\partial_{2}A_{1})|\phi|^{2}-2(\partial_{1}\phi_{1}-A_{1}\phi_{2})(\partial_{2}\phi_{2}+A_{2}\phi_{1})+2(\partial_{1}\phi_{2}+A_{1}\phi_{1})(\partial_{2}\phi_{1}-A_{2}\phi_{2})$
$=\underbrace{(-\partial_{1}A_{2}\cdot\phi^{2}_{1}-2A_{2}\phi_{2}\cdot\partial_{1}\phi_{1})}_{\text{$-\partial_{1}(A_{2}\phi^{2}_{1})$}}+\underbrace{(-\partial_{1}A_{2}\cdot\phi^{2}_{2}-2A_{2}\phi_{2}\cdot\partial_{1}\phi_{2})}_{\text{$-\partial_{1}(A_{2}\phi^{2}_{2})$}}+$
$+\underbrace{(\partial_{2}A_{1}\cdot\phi^{2}_{1}+2A_{1}\phi_{1}\cdot\partial_{1}\phi_{2})}_{\text{$\partial_{2}(A_{1}\phi^{2}_{1})$}}+\underbrace{(\partial_{2}A_{1}\cdot\phi^{2}_{2}+2A_{1}\phi_{2}\cdot\partial_{2}\phi_{2})}_{\text{$\partial_{2}(A_{1}\phi^{2}_{2})$}}+$
$+\underbrace{2(-\partial_{1}\phi_{1}\cdot\partial_{2}\phi_{2}+\partial_{1}\phi_{2}\cdot\partial_{2}\phi_{1})}_{\text{$-2\det(d\phi)$}}.$
Note that
$d(\phi_{1}d\phi_{2}-\phi_{2}d\phi_{1})$ $=2d\phi_{1}\wedge d\phi_{2}=2\det(d\phi)dx_{1}\wedge dx_{2},$
$d(iA|\phi|^{2})=-d(A_{2}|\phi|^{2}dx_{2}+A_{1}|\phi|^{2}dx_{1})$ $=(-\partial_{1}(A_{2}|\phi|^{2})+\partial_{2}(A_{1}|\phi|^{2}))dx_{1}\wedge dx_{2}.$
Combine all of the above and we obtain the new formula for $\mathscr{Y}(A,\phi)$ as claimed. ∎
Proposition 3.2.
Let $(A,\phi)$ such that $A$ is a continuous connection and $\phi\in C^{1}$. Suppose we have
$\lim_{R\to\infty}\sup_{|x|=R}|1-|\phi|^{2}|=0,\quad|x|^{1+\delta}|\nabla_{A}\phi|\leq c_{0},$
for some $\delta,c_{0}>0$. Then, $\mathscr{Y}(A,\phi)\geq\dfrac{1}{2}\displaystyle\iint_{\mathbf{R}^{2}}F(A)$. Equality happens if and only if $(A,\phi)$ satisfies
(3.2) $\begin{cases}\dfrac{\partial\phi}{\partial\overline{z}}-\dfrac{i}{2}(A_{1}+iA_{2})\phi=0,\\ \dfrac{i}{2}F(A)=(1-|\phi|^{2})dz\wedge d\overline{z},\end{cases}$
where $A=iA_{1}dx_{1}+iA_{2}dx_{2}$.
Proof.
Note that $\phi_{2}(d\phi_{1}-iA\phi_{2})-\phi_{1}(d\phi_{2}+iA\phi_{1})$ can be re-written as
$\left\langle\begin{bmatrix}d&-iA\\ iA&d\end{bmatrix}\,\begin{bmatrix}\phi_{1}\\ \phi_{2}\end{bmatrix},\begin{bmatrix}\phi_{2}\\ -\phi_{1}\end{bmatrix}\right\rangle=\Re\langle\nabla_{A}\phi,-i\phi\rangle,$
[…]
$\displaystyle\iint_{\mathbf{R}^{2}}d(\phi_{2}(d\phi_{1}-iA\phi_{2})-\phi_{1}(d\phi_{2}+iA\phi_{1}))$
$=\lim_{R\to\infty}\iint_{|x|\leq R}d(\phi_{2}(d\phi_{1}-iA\phi_{2})-\phi_{1}(d\phi_{2}+iA\phi_{1}))$
$=\lim_{R\to\infty}\int_{|x|=R}\Re\langle\nabla_{A}\phi,-i\phi\rangle$
$\leq\lim_{R\to\infty}\int_{|x|=R}|\nabla_{A}\phi|\cdot|\phi|.$
Let $c_{R}$ denote the supremum of $|1-|\phi|^{2}|$ on $|x|=R$. Then $|\phi|\leq\sqrt{1+c_{R}}$ for any $|x|=R$. By the hypothesis of the proposition, the last integral on the right-hand side of the above inequality can be estimated further by
$\lim_{R\to\infty}\int_{|x|=R}|\nabla_{A}\phi|\cdot|\phi|\leq\lim_{R\to\infty}\int_{|x|=R}\dfrac{c_{0}}{R^{1+\delta}}\cdot\sqrt{1+c_{R}}=\lim_{R\to\infty}\dfrac{2\pi c_{0}}{R^{\delta}}\cdot\sqrt{1+c_{R}}=0.$
Therefore, the integral of $d(\phi_{2}(d\phi_{1}-iA\phi_{2})-\phi_{1}(d\phi_{2}+iA\phi_{1}))$ over $\mathbf{R}^{2}$ is equal to zero. Hence, by Lemma 3.2, we immediately get the estimate of $\mathscr{Y}$ as claimed. The statement of equality can be checked directly via calculations (also, see [MR0614447], Ch.3). ∎
One can say more.
Theorem 3.3 (Proposition 3.5, Theorem 1.1 [MR0614447]).
Suppose $(A,\phi)$ satisfies the condition of Proposition 3.2 and $F(A)\in L^{1}$. Then
  1. (1)
    $\dfrac{1}{2\pi}\iint_{\mathbf{R}^{2}}F(A)=N$, where $N$ is an integer.
  2. (2)
    If $N\geq 0$, given a set $\{z_{j}\}_{j=1,\cdots N}$ in $\mathbf{C}$, there is a finite action solution $(A,\phi)$ (i.e., $\mathscr{Y}(A,\phi)<\infty$) of (3.2) such that
    1. (a)
      $(A,\phi)$ is globally smooth.
    2. (b)
      The zeroes of $\phi$ are $\{z_{j}\}$. And as $z\to z_{j}$, we have $\phi(z,\overline{z})\sim c_{j}(z-z_{j})^{n_{j}}$, where $c_{j}\neq 0$ and $n_{j}$ is the multiplicity of $z_{j}$.
Remark 3.4.
(3.2) is called a vortex equation. Solutions of the vortex equation (3.2) are also solutions of the Euler-Lagrange equation of $\mathscr{Y}$. Theorem 1.2 in [MR0614447] shows that the only finite action critical point of $\mathscr{Y}$ is a solution of (3.2) in the form in Theorem 3.3. Note the first equation in (3.2) is a Vekua equation (cf. (2.3)).
Remark 3.5.
Both Lemma 3.1 and Proposition 3.2 are stated in [MR0614447], but a proof of them was not given in detail. We present it here for the sake of self-containment.
There is a symmetry of $\mathcal{C}$ that makes $\mathscr{Y}$ invariant. The symmetry is given by the gauge group $\mathcal{G}=Maps(\mathbf{R}^{2},U(1))$, where the action $\mathcal{G}\curvearrowright\mathcal{C}$ is given by
$(\sigma,(A,\phi))\mapsto(A+\sigma d\sigma^{-1},\sigma\cdot\phi).$
By direct calculations, it is not difficult to see that $\mathscr{Y}(\sigma\cdot(A,\phi))=\mathscr{Y}(A,\phi)$. Thus, $\mathscr{Y}$ descends to a function (also denoted by the same name when the context is clear) $\mathscr{Y}:\mathcal{C}/\mathcal{G}\to\mathbf{R}$. As a result, solutions of the vortex equation (3.2) are also $\mathcal{G}$-invariant.
3.2. Seiberg-Witten Gauge Theory
There is another variant of the vortex equation (3.2) that is derived from a slightly different perspective. For that, we make a detour to dimension four and briefly discuss a gauge theoretic equation called the Seiberg-Witten equations. The Seiberg-Witten equations can be defined on any $4$-manifold. However, following the theme of the previous subsection, we mainly focus on its formulation in the Euclidean space $\mathbf{R}^{4}$.
Consider the standard flat metric on $\mathbf{R}^{4}$, $\{x_{0},\cdots,x_{3}\}$ are the coordinates, $\{e_{0},\cdots,e_{3}\}$ are the standard orthonormal basis of its tangent bundle $T\mathbf{R}^{4}=\mathbf{R}^{4}\times\mathbf{R}^{4}$, $\{dx_{0},\cdots,dx_{3}\}$ are the dual bases for $T^{*}\mathbf{R}^{4}$. Fix the constant $spin^{c}$ structure $\rho:\mathbf{H}=\mathbf{R}^{4}\to\mathbf{C}^{4\times 4}$ (see [MR3012377, MR1367507]) defined by
$\rho(v)=\begin{bmatrix}0&v\\ -\overline{v}^{t}&0\end{bmatrix},\,\,\,\,v=\begin{bmatrix}a+bi&c+di\\ -c+di&a-bi\end{bmatrix}.$
So we identify $e_{0}=Id,e_{1}=I,e_{2}=J,$ and $e_{3}=K$ with
$I=\begin{bmatrix}i&0\\ 0&-i\end{bmatrix},\,\,\,J=\begin{bmatrix}0&1\\ -1&0\end{bmatrix},\,\,\,K=\begin{bmatrix}0&i\\ i&0\end{bmatrix}.$
Let $S^{+}$ denote the above $spin^{c}$ structure and $L_{\rho}=\mathbf{R}^{4}\times\mathbf{C}$ be the associated line bundle. Consider the $spin^{c}$ connection $\nabla=\nabla_{A}$ given by
$\nabla_{j}\,\psi=\dfrac{\partial\psi}{\partial x_{j}}+iA_{j}\,\psi,\,\,\,j=0,\cdots,3,$
where $A_{j}:\mathbf{R}^{4}\to\mathbf{R}$ and $\psi:\mathbf{R}^{4}\to\mathbf{C}$. The associated connection on $L_{\rho}$ is given by $A=iA_{0}\,dx_{0}+iA_{1}\,dx_{1}+iA_{2}\,dx_{2}+iA_{3}\,dx_{3}$. Note that $\psi:\mathbf{R}^{4}\to\mathbf{C}$ is called a spinor and $\nabla_{A}$ is called a spinor connection. We denote $\mathcal{C}(S^{+})=i\Omega^{1}(\mathbf{R}^{4})\times C^{\infty}(\mathbf{R}^{4},\mathbf{C}^{2})$ by the configuration of the $spin^{c}$ structure $S^{+}$.
Definition 3.6.
Given a spinor connection $A$ on $\mathbf{R}^{4}$. The Dirac operator $D^{+}_{A}$ defined on $C^{\infty}(\mathbf{R}^{4},\mathbf{C}^{2})$ is an elliptic first order operator given by
$D^{+}_{A}\psi=-\nabla_{0}\psi+I\nabla_{1}\psi+J\nabla_{2}\psi+K\nabla_{3}\psi.$
In a general Euclidean space $\mathbf{R}^{n}$ with the standard flat metric, there is the Hodge $\star$-operator that takes a $p$-form to a $(n-p)$-form. It is defined as follows. Let $\omega$ be a $p$-form on $\mathbf{R}^{n}$ written in Einstein summation notation as
$\omega=\omega_{j_{1}\cdots j_{p}}dx_{j_{1}}\wedge\cdots\wedge dx_{j_{p}}.$
Then
$\star\omega=\dfrac{1}{p!}\epsilon^{k_{1}\cdots k_{p}j_{1}\cdots j_{n-p}}\omega_{k_{1}\cdots k_{p}}dx_{j_{1}}\wedge\cdots\wedge dx_{j_{n-p}},$
where $\epsilon^{k_{1}\cdots k_{p}}$ is the totally anti-symmetric tensor and $\epsilon^{1\cdots n}=1$. In dimension four, $\star$-operator turns a $2$-form to another $2$-form. Since $\star^{2}=1$, its eigenvalues are ±$\pm 1$. We say that a two form $\omega$ is (anti) self-dual if and only if $\star\omega=\pm\omega$. Any two form $\omega$ can be written as a sum of a self-dual form and an anti-self-dual form,
$\omega=\underbrace{\omega^{+}}_{\text{$\dfrac{1}{2}(\omega+\star\omega)$}}+\underbrace{\omega^{-}}_{\text{$\dfrac{1}{2}(\omega-\star\omega)$}}.$
Another special feature of dimension four is that $\rho$ is also an isometry between the space of purely imaginary self-dual $2$-forms on $\mathbf{R}^{4}$ and the space of all self-adjoint traceless endomorphisms of $\mathbf{C}^{2}$. In particular, the self-dual part $F^{+}(A)$ of the curvature of $A$ can be written as
$F^{+}(A)=i(F_{01}+F_{23})I+i(F_{02}+F_{31})J+i(F_{03}+F_{12})K,\,\,\text{ where }F_{jk}=\dfrac{\partial A_{k}}{\partial x_{j}}-\dfrac{\partial A_{j}}{\partial x_{k}}.$
There is another way to obtain a self-adjoint traceless endomorphism of $\mathbf{C}^{2}$ from a spinor $\psi$. We define $\mu(\psi)$ as a linear combination of $I,J,K$ in the following way
$\mu(\psi)=\dfrac{1}{2}(\psi^{*}I\psi)I+\dfrac{1}{2}(\psi^{*}J\psi)J+\dfrac{1}{2}(\psi^{*}K\psi)K.$
Here $\psi^{*}$ denotes the conjugate transpose of $\psi$ when we view it as a column vector in $\mathbf{C}^{2}$. Having set these up, we are ready to write down the Seiberg-Witten equations in $\mathbf{R}^{4}$.
Definition 3.7.
The Seiberg-Witten equations on $\mathbf{R}^{4}$ is a system of non-linear elliptic PDEs that look for the unknown $(A,\psi)\in\mathcal{C}(S^{+})$ satisfying
(3.3) $\begin{cases}D^{+}_{A}\psi=0\\ i(F_{01}+F_{23})=\dfrac{1}{2}\psi^{*}I\psi\\ i(F_{02}+F_{31})=\dfrac{1}{2}\psi^{*}J\psi\\ i(F_{03}+F_{12})=\dfrac{1}{2}\psi^{*}K\psi.\end{cases}$
Remark 3.8.
If we write $\psi=\begin{bmatrix}\psi_{1}&\psi_{2}\end{bmatrix}^{T}$, where $\psi_{j}:\mathbf{R}^{4}\to\mathbf{C}$, then note that $\psi^{*}I\psi=i(|\psi_{1}|^{2}-|\psi_{2}|^{2})$, $\psi^{*}J\psi=2i\Im(\overline{\psi_{1}}\psi_{2})$, and $\psi^{*}K\psi=2i\Re(\overline{\psi_{1}}\psi_{2})$. These expressions are all homogeneous polynomials in $\psi_{1},\psi_{2}$ variables of degree $2$.
3.3. Dimensional Reduction
The following dimensional reduction of (3.3) from $\mathbf{R}^{4}$ to $\mathbf{R}^{2}$ is in the spirit of Hitchin’s work on self-dual Yang-Mills equations on Riemann surfaces [MR0887284]. It has been done also in Dey’s thesis [MR2698453, MR1952133]. We will assume that $(A,\psi)\in\mathcal{C}(S^{+})$ is invariant in the $x_{2},x_{3}$-coordinate. Let’s first take a look at the curvature equations in (3.3). In this setup, we have
(3.4) $\begin{cases}iF_{01}=\dfrac{i}{2}(|\psi_{1}|^{2}-|\psi_{2}|^{2})\\ i\left(\dfrac{\partial A_{2}}{\partial x_{0}}-\dfrac{\partial A_{3}}{\partial x_{1}}\right)=i\Im(\overline{\psi_{1}}\psi_{2})\\ i\left(\dfrac{\partial A_{3}}{\partial x_{0}}+\dfrac{\partial A_{2}}{\partial x_{1}}\right)=i\Re(\overline{\psi_{1}}\psi_{2}).\end{cases}$
If we denote $\phi_{0}=iA_{2}$ and $\phi_{1}=iA_{3}$, then we can rewrite (3.4) as
(3.5) $\begin{cases}iF_{01}=\dfrac{i}{2}(|\psi_{1}|^{2}-|\psi_{2}|^{2})\\ \dfrac{\partial\phi_{0}}{\partial x_{0}}-\dfrac{\partial\phi_{1}}{\partial x_{1}}=i\Im(\overline{\psi_{1}}\psi_{2})\\ \dfrac{\partial\phi_{0}}{\partial x_{1}}+\dfrac{\partial\phi_{1}}{\partial x_{0}}=i\Re(\overline{\psi_{1}}\psi_{2}).\end{cases}$
Introduce the complex coordinate $z=x_{0}+x_{1}\,i$ and note that $\partial/\partial z=(\partial/\partial x_{0}-i\partial/\partial x_{1})/2$ and $\partial/\partial\overline{z}=(\partial/\partial x_{0}+i\partial/\partial x_{1})/2$, combine the last two equations of (3.5) to obtain
(3.6) $\dfrac{\partial\phi}{\partial\overline{z}}=-\dfrac{1}{2}\psi_{1}\overline{\psi_{2}},\,\,\text{where }\phi=\phi_{0}+i\phi_{1}.$
If we let $\Phi=\phi dz-\overline{\phi}d\overline{z}$, then we can re-write (3.4) in forms as
(3.7) $\begin{cases}2\overline{\partial}\,\Phi=-\dfrac{1}{2}\psi_{1}\overline{\psi_{2}}\,dz\wedge d\overline{z}\\ F(A)=\dfrac{i}{2}(|\psi_{1}|^{2}-|\psi_{2}|^{2})\,dz\wedge d\overline{z}\end{cases}$
Moving on to the Dirac equation of (3.3), in our setup, we view $\psi=\begin{bmatrix}\psi_{1}&\psi_{2}\end{bmatrix}^{T}$, where $\psi_{1},\psi_{2}:\mathbf{C}\to\mathbf{C}$. Then $D^{+}_{A}\psi=0$ can be re-written as follows
$\begin{bmatrix}\dfrac{\partial}{\partial x_{0}}\psi_{1}+iA_{0}\psi_{1}\\ \dfrac{\partial}{\partial x_{0}}\psi_{2}+iA_{0}\psi_{2}\end{bmatrix}-\begin{bmatrix}i\dfrac{\partial}{\partial x_{1}}\psi_{1}-A_{1}\psi_{1}\\ -i\dfrac{\partial}{\partial x_{1}}\psi_{2}+A_{1}\psi_{2}\end{bmatrix}-\begin{bmatrix}iA_{2}\psi_{2}\\ -iA_{2}\psi_{1}\end{bmatrix}-\begin{bmatrix}-A_{3}\psi_{2}\\ -A_{3}\psi_{1}\end{bmatrix}=0$
$\Rightarrow\begin{bmatrix}\dfrac{\partial}{\partial x_{0}}\psi_{1}+iA_{0}\psi_{1}-i\dfrac{\partial}{\partial x_{1}}\psi_{1}+A_{1}\psi_{1}-iA_{2}\psi_{2}+A_{3}\psi_{2}\\ \dfrac{\partial}{\partial x_{0}}\psi_{2}+iA_{0}\psi_{2}+i\dfrac{\partial}{\partial x_{1}}-A_{1}\psi_{2}+iA_{2}\psi_{1}+A_{3}\psi_{1}\end{bmatrix}=0$
(3.8) $\Rightarrow\begin{bmatrix}2\dfrac{\partial}{\partial z}+i(A_{0}-iA_{1})&-\phi\\ -\overline{\phi}&2\dfrac{\partial}{\partial\overline{z}}+i(A_{0}+iA_{1})\end{bmatrix}\cdot\begin{bmatrix}\psi_{1}\\ \psi_{2}\end{bmatrix}=0$
If we let $A^{1,0}=i(A_{0}-iA_{1})/2\,dz$ and $A^{0,1}=i(A_{0}+iA_{1})\,d\overline{z}$ so that $A=iA_{0}dx_{0}+iA_{1}dx_{1}=A^{1,0}+A^{0,1}$, then we can re-write (3.8) one more time as:
(3.9) $\begin{bmatrix}\dfrac{1}{2}\Phi^{0,1}&-(\overline{\partial}+A^{0,1})\\ \partial+A^{1,0}&-\dfrac{1}{2}\Phi^{1,0}\end{bmatrix}\cdot\begin{bmatrix}\psi_{1}\\ \psi_{2}\end{bmatrix}=0.$
(3.7) and (3.9) together gives us the dimensional reduction of the Seiberg-Witten equations (3.3) over $\mathbf{R}^{2}$
(3.10) $\begin{cases}2\overline{\partial}\,\Phi=-\dfrac{1}{2}\psi_{1}\overline{\psi_{2}}\,dz\wedge d\overline{z}\\ F(A)=\dfrac{i}{2}(|\psi_{1}|^{2}-|\psi_{2}|^{2})\,dz\wedge d\overline{z}\\ \begin{bmatrix}-\dfrac{1}{2}\Phi^{0,1}&-(\overline{\partial}+A^{0,1})\\ \partial+A^{1,0}&-\dfrac{1}{2}\Phi^{1,0}\end{bmatrix}\cdot\begin{bmatrix}\psi_{1}\\ \psi_{2}\end{bmatrix}=0.\end{cases}$
Solutions of (3.10) are $(A,\psi_{1},\psi_{2},\Phi)$, where $A\in i\Omega^{1}(\mathbf{R}^{2})$, $\Phi\in\Omega^{1,1}(\mathbf{R}^{2})$, and $\psi_{1},\psi_{2}:\mathbf{C}\to\mathbf{C}$. One should think of $A$ as some associated connection to the trivial complex line bundle over $\mathbf{R}^{2}$, and $\Phi$ is a Higgs field. Without the Higgs field, there is another variant of the vortex equations given by
(3.11) $\begin{cases}F(A)=\dfrac{i}{2}(|\psi_{1}|^{2}-|\psi_{2}|^{2})\,dz\wedge d\overline{z}\\ \begin{bmatrix}0&-(\overline{\partial}+A^{0,1})\\ \partial+A^{1,0}&0\end{bmatrix}\cdot\begin{bmatrix}\psi_{1}\\ \psi_{2}\end{bmatrix}=0.\end{cases}$
The "non-forms" version of (3.10) will be read as
(3.12) $\begin{cases}\dfrac{\partial\phi}{\partial\overline{z}}=-\dfrac{1}{2}\psi_{1}\overline{\psi_{2}},\\ i\left(\dfrac{\partial A_{1}}{\partial x_{0}}-\dfrac{\partial A_{0}}{\partial x_{1}}\right)=\dfrac{i}{2}(|\psi_{1}|^{2}-|\psi_{2}|^{2})\\ 2\dfrac{\partial\psi_{2}}{\partial\overline{z}}+i(A_{0}+iA_{1})\psi_{2}-\overline{\phi}\psi_{1}=0\\ 2\dfrac{\partial\psi_{1}}{\partial z}+i(A_{0}-iA_{1})\psi_{1}-\phi\psi_{2}=0\end{cases}$
The gauge group action on a configuration is defined similarly as at the end of Subsection 3.1. The space of solutions of these equations quotient out by $\mathcal{G}$ are called the moduli spaces of the Seiberg-Witten vortex equations.
Proposition 3.9.
There are non-trivial solutions to (3.11). Explicitly, for any $(c_{1},c_{2})\in\mathbf{C}^{*}\times\mathbf{C}^{*}$, $(A_{0},A_{1},\psi_{1},\psi_{2})=(-2c_{2},0,\pm c_{1}e^{ic_{2}(z+\overline{z})},c_{1}e^{ic_{2}(z+\overline{z})})$ is a solution of (3.11).
Proof.
The proof is just a direct calculation. ∎
Note that the solutions above have no zeroes in $\mathbf{C}$. Suppose now we would like to use Proposition 3.9 as a building block for solutions with prescribed zeroes. The process would start as follows. If we wish to use $\psi_{2}$ as in Proposition 3.9 and obtain another solution, by the Vekua representation result 2.4, we have to rescale $\psi_{2}$ by a holomorphic function $h_{2}$. Keeping $A_{0},A_{1}$ the same as in the above proposition, then $(A_{0},A_{1},\psi_{2})=(-2c_{2},0,c_{1}h_{2}e^{ic_{2}(z+\overline{z})})$ is a solution of $02\partial_{\overline{z}}\psi_{2}+i(A_{0}+iA_{1})\psi_{2}=$ 0 if and only if $\partial_{\overline{z}}h_{2}=0$. Similarly, we also have to rescale $\psi_{1}$ by an anti-holomorphic $h_{1}$. For $(A_{0},A_{1},\psi_{1})=(-2c_{2},0,c_{1}h_{1}e^{ic_{2}(z+\overline{z})})$ to be a solution of $02\partial_{z}\psi_{1}+i(A_{0}-iA_{1})\psi_{1}=$ 0, we need $\partial_{z}(h_{1})=0$. Now, for $(A_{0},A_{1},\psi_{1},\psi_{2})=(-2c_{2},0,c_{1}h_{1}e^{ic_{2}(z+\overline{z})},c_{1}h_{2}e^{ic_{2}(z+\overline{z})})$ to be a solution of the curvature equation, we need $|h_{1}|=|h_{2}|$. To sum up, we have
Corollary 3.10.
$(A_{0},A_{1},\psi_{1},\psi_{2})=(-2c_{2},0,c_{1}h_{1}e^{ic_{2}(z+\overline{z})},c_{1}h_{2}e^{ic_{2}(z+\overline{z})})$ is another solution of (3.11) if and only if
(3.13) $\begin{cases}\partial_{\overline{z}}h_{2}=0,\\ \partial_{z}(h_{1})=0,\\ |h_{1}|=|h_{2}|.\end{cases}$
Since $\overline{h_{1}},h_{2}$ are holomorphic functions that share the same modulus, there is a constant $\lambda\in S^{1}$ such that $\overline{h_{1}}=\lambda h_{2}$. As a result, this gives us a recipe to yield other solutions. Let $\{z_{1},\cdots,z_{k}\}$ be some points in $\mathbf{C}$. Let $\theta$ be any real number and consider
$h_{2}(z)=(z-z_{1})^{n_{1}}\cdots(z-z_{k})^{n_{k}},\quad h_{1}(z)=e^{i\theta}(\overline{z}-\overline{z_{1}})^{n_{1}}\cdots(\overline{z}-\overline{z_{k}})^{n_{k}}.$
Then $(A_{0},A_{1},\psi_{1},\psi_{2})$ given by
(3.14) $(-2c_{2},0,c_{1}e^{i\theta}(\overline{z}-\overline{z_{1}})^{n_{1}}\cdots(\overline{z}-\overline{z_{k}})^{n_{k}}e^{ic_{2}(z+\overline{z})},c_{1}(z-z_{1})^{n_{1}}\cdots(z-z_{k})^{n_{k}}e^{ic_{2}(z+\overline{z})}).$
is always a solution of (3.11). As a final remark, such a solution always has a flat connection by construction. Later, we will show the existence of a different kind of solution to (3.11) (where the connection is not necessarily flat) with a finite set of zeroes via a calculus of variations method. Denote by $Vor_{\mathfrak{p}}$ the moduli space of solutions to (3.11) of type (3.14). We summarize the discussion above in the form of the following theorem.
Theorem 3.11.
There is a surjective map $\eta_{\mathfrak{p}}:Vor_{\mathfrak{p}}(\mathbf{C})\to\bigcup_{n\in\mathbf{N}}Sym^{n}(\mathbf{C})$.
4. Existence Of Solutions With Exponential Decay
4.1. A Kazdan-Warner Equation
In this section, we shall prove that the Seiberg-Witten vortex equations without Higgs field (3.11) has a solution (where the connection is not necessarily flat) under appropriate assumptions about the behavior at the point at infinity. Firstly, we note that the "non-forms" version of (3.11) can be written as follows
(4.1) $\begin{cases}2\dfrac{\partial\psi_{2}}{\partial\overline{z}}+i(A_{0}+iA_{1})\psi_{2}=0,\\ 2\dfrac{\partial\psi_{1}}{\partial z}+i(A_{0}-iA_{1})\psi_{1}=0,\\ i\left(\dfrac{\partial A_{1}}{\partial x_{0}}-\dfrac{\partial A_{0}}{\partial x_{1}}\right)=\dfrac{i}{2}(|\psi_{1}|^{2}-|\psi_{2}|^{2}).\end{cases}$
We let $\alpha=\dfrac{1}{2}(A_{0}-iA_{1})$. Then by direct calculations, we can re-write (4.1) as
(4.2) $\begin{cases}\partial_{\overline{z}}\psi_{2}+i\overline{\alpha}\psi_{2}=0,\\ \partial_{z}\psi_{1}+i\alpha\psi_{1}=0,\\ i(\partial_{z}\overline{\alpha}-\partial_{\overline{z}}\alpha)=\dfrac{1}{2}(|\psi_{1}|^{2}-|\psi_{2}|^{2}).\end{cases}$
In the first equation of (4.2), we solve for $\alpha$ to obtain
(4.3) $i\overline{\alpha}=-\partial_{\overline{z}}\log\psi_{2},\quad\psi_{2}\neq 0.$
By substituting the expression of $\alpha$ into the second equation of (4.2) and conjugate, we have
(4.4) $\partial_{\overline{z}}\log(\overline{\psi_{1}}\psi_{2})=0,\quad\psi_{1},\psi_{2}\neq 0.$
We would like to comment that (4.4) can only make sense at the points in $\mathbf{C}$ where $\psi_{1}$ and $\psi_{2}$ do not vanish. Denote $S_{n}=\{z_{1},\cdots,z_{n}\}$ by the set of all zeroes of $\psi_{1}$ and $\psi_{2}$. Let $\theta_{j}$ be a multivalued function such that $\psi_{j}=\exp\left(\dfrac{1}{2}|\psi_{j}|^{2}+\dfrac{i}{2}\theta_{j}\right)$. Then $\log(\overline{\psi_{1}}\psi_{2})=\dfrac{1}{2}(|\psi_{1}|^{2}+|\psi_{2}|^{2}-i(\theta_{1}-\theta_{2}))$. For now, assume $\psi_{1},\psi_{2}$ are such that the only solution to (4.4) are constants holomorphic function on $\mathbf{C}$. Then for each constant $C\in\mathbf{C}$, we must have $|\psi_{1}|^{2}+|\psi_{2}|^{2}-i(\theta_{1}-\theta_{2})=C$. If we write $C=C_{1}+iC_{2}$, then $|\psi_{1}|^{2}=C_{1}-|\psi_{2}|^{2}$. Given this, as we substitute the expression for $\alpha$ in (4.3) into the remaining equation of the system (4.2), we obtain
(4.5) $-\dfrac{1}{4}\Delta\log|\psi_{2}|^{2}=\dfrac{1}{2}(C_{1}-2|\psi_{2}|^{2})\Leftrightarrow-\Delta u+4e^{u}-2C_{1}=0.$
Here $u=\log|\psi_{2}|^{2}$. Note that (4.5) only makes sense on $\mathbf{C}\setminus S_{n}$. Let $\delta$ be the Dirac measure. To account for the singularities of $u$, we consider the following equation
(4.6) $-\Delta u+4e^{u}-2C_{1}=-4\pi\displaystyle\sum_{k=1}^{n}\delta(z-z_{k}).$
Therefore, the existence of solutions of (4.1) is equivalent to the existence of solutions of (4.6). We now address the condition on $\psi_{1},\psi_{2}$ such that $\log(\overline{\psi_{1}}\psi_{2})$ satisfying the Cauchy-Riemann equation can only be constant. The easiest way for such a condition to be achieved is to look for solutions of (4.1), where $\psi_{1}$ and $\psi_{2}$ have the same argument, i.e., $\psi_{1}=\lambda\psi_{2}$, with $\lambda$ being a $\mathbf{R}$-valued function. If those are the solutions that we look for, then from the equation $\partial_{\overline{z}}\log(\overline{\psi_{1}}\psi_{2})=0$, we must have $\partial_{\overline{z}}(|\psi_{1}|^{2}+|\psi_{2}|^{2})=0$. As a result, $|\psi_{1}|^{2}+|\psi_{2}|^{2}=C$. Note that $C\geq 0$. If $C=0$, then we $A$ must be a closed $1$-form on $\mathbf{R}^{2}$. But on $\mathbf{R}^{2}$, every closed $1$-form is also exact. Thus, $A=df$, where $f$ is any smooth $i\mathbf{R}$-valued function on $\mathbf{R}^{2}$. In this case, we obtain a family of solutions of (4.1) $\{(df,0,0)\}_{f\in C^{\infty}(\mathbf{R}^{2},i\mathbf{R})}$. Now, if $C>0$ and denote $u=\log|\psi_{2}/\sqrt{C}|^{2}$, then away from $S_{n}$, we have $-\Delta u+4Ce^{u}-2C=0$. Thus, on the entire $\mathbf{C}$, $u$ is a solution of
(4.7) $-\Delta u+4Ce^{u}-2C=-4\pi\displaystyle\sum_{k=1}^{n}\delta(z-z_{k}),$
with the condition that $u\leq 0$ and $\lim_{|z|\to\infty}u=0$. The second condition follows from the assumption that $\lim_{R\to\infty}\sup_{|z|=R}|1-|\psi_{2}|^{2}/C|=0$. The condition is similar to the one that appears in Proposition 3.2. However, unlike the vortex equations considered by Taubes, we are unaware of whether (4.1) is a critical condition for any "useful" energy functional. Regardless, we artificially impose this condition to ensure the existence of non-trivial solutions to (4.7). By the way, as a remark, (4.7) is thought of as a Kazdan-Warner equation.
The main theorem of this section we will prove is the following.
Theorem 4.1.
Let $C$ be any positive number. There exists a smooth solution $(A_{0},A_{1},\psi_{1},\psi_{2})$ of (4.1) with the property (E) that $\psi_{1},\psi_{2}$ have the same argument and that for every $\epsilon>0$, there exists a $c>0$ such that $C-|\psi_{2}|^{2}\leq c\cdot\exp(-(1-\epsilon)|x|)$.
Remark 4.2.
Denote by $Vor_{\mathfrak{e}}(\mathbf{C})$ the moduli space of solutions of (4.1) with property (E). There is well-defined map from $\eta_{\mathfrak{e}}:Vor_{\mathfrak{e}}(\mathbf{C})\to\bigcup_{n\in\mathbf{N}}Sym^{n}(\mathbf{C})$. It is interesting to investigate more about the properties of such map $\eta_{\mathfrak{e}}$. We will address this direction elsewhere.
Remark 4.3.
Note that the construction of the solutions at the end of the previous section is not of property (E). In fact, they all have polynomial growth. Recall that the collection of the gauge equivalence classes of such solutions is denoted by $Vor_{\mathfrak{p}}(\mathbf{C})$. Similarly, there is an obvious map $\eta_{\mathfrak{p}}:Vor_{\mathfrak{p}}(\mathbf{C})\to\bigcup_{n\in\mathbf{N}}Sym^{n}(\mathbf{C})$. It is not difficult to see that $\eta_{\mathfrak{p}}$ is surjective (cf. Theorem 3.11).
The first part of Theorem 4.1 is an immediate consequence of the following theorem.
Theorem 4.4.
For $C>0$, the equation $-\Delta u+4Ce^{u}-2C=-4\pi\displaystyle\sum_{k=1}^{n}\delta(z-z_{k})$ has a real analytic solution with the condition that $u\leq 0$ and $u\to 0$ in a certain appropriate sense.
The proof of Theorem 4.4 will be given in the following subsections. Next, we discuss some aspects that go into the proof. The strategy we employ here is similar to that of Taubes’ in [MR0614447] and will be divided into several steps. (cf. 4.1)
Step 1.
Simplify (4.7). Note that the equation (4.7) has a term involving the Dirac measure. To simplify the analysis, we re-write (4.7) by shifting $u$ with a Green-like fundamental solution viewed as a distribution $u_{0}$. The new equivalent equation would have no term involving the Dirac measure (cf. Subsection 4.1). ∎
Step 2.
Realize the simplified version of (4.7) as an Euler-Lagrange equations for certain functional $\mathscr{A}$ defined on $C^{\infty}_{c}(\mathbf{R}^{2},\mathbf{R})$ (cf. Subsection 4.1). ∎
Step 3.
Show that $\mathscr{A}$ is well-defined for a certain Banach space $\mathcal{B}$ completion of $C^{\infty}_{c}(\mathbf{R}^{2},\mathbf{R})$. Typically, $\mathcal{B}$ is taken to be a certain Sobolev completion of $C^{\infty}_{c}(\mathbf{R}^{2},\mathbf{R})$ (cf. Subsection 4.2). ∎
Step 4.
Show that $\mathscr{A}:\mathcal{B}\to\mathbf{R}$ achieves a minimum at a critical point (cf. Subsection 4.3). ∎
Step 5.
Recover analyticity of a critical point of $\mathscr{A}:\mathcal{B}\to\mathbf{R}$ by appealing to elliptic theory (cf. Subsection 4.3). ∎
To end this subsection, we look at Steps 1 and 2 laid out above.
Lemma 4.5.
Let $\mu$ be a real number such that $\mu>2n/C$. Denote the distribution $u_{0}$ by $u_{0}=-\sum_{k=1}^{n}\log(1+\mu|z-z_{k}|^{-2})$. Then
$\Delta u_{0}=-4\displaystyle\sum_{k=1}^{n}\dfrac{\mu}{(|z-z_{k}|^{2}+\mu)^{2}}-4\pi\sum_{k=1}^{n}\delta(z-z_{k}).$
Proof.
The proof is via direct calculations. Firstly, away from the singular set $S_{n}$, we have
$\Delta u_{0}$ $=-\sum_{k=1}^{n}4\partial_{z}\partial_{\overline{z}}\log(1+\mu|z-z_{k}|^{-2})$
$=-\sum_{k=1}^{n}4\partial_{z}\left\{\dfrac{1}{1+\mu|z-z_{k}|^{-2}}\cdot\partial_{\overline{z}}(1+\mu|z-z_{k}|^{-2})\right\}$
$=-\sum_{k=1}^{n}4\mu\cdot\partial_{z}\left\{\dfrac{-(z-z_{k})}{1+\mu|z-z_{k}|^{-2}}\cdot\dfrac{1}{|z-z_{k}|^{4}}\right\}$
$=\sum_{k=1}^{n}4\mu\cdot\partial_{z}\left\{\dfrac{z-z_{k}}{|z-z_{k}|^{4}+\mu|z-z_{k}|^{2}}\right\}$
$=\sum_{k=1}^{n}4\mu\cdot\dfrac{|z-z_{k}|^{4}+\mu|z-z_{k}|^{2}-(z-z_{k})(2(\overline{z}-\overline{z_{k}})^{2}(z-z_{k})+\mu(\overline{z}-\overline{z_{k}}))}{(|z-z_{k}|^{4}+\mu|z-z_{k}|^{2})^{2}}$
$=-4\sum_{k=1}^{n}\dfrac{\mu|z-z_{k}|^{4}}{(|z-z_{k}|^{4}+\mu|z-z_{k}|^{2})^{2}}=-4\sum_{k=1}^{n}\dfrac{\mu}{(|z-z_{k}|^{2}+\mu)^{2}}.$
As a result, viewing $u_{0}$ as a distribution, near each singularity we obtain the formula for $\Delta u_{0}$ as claimed. ∎
Note that the function $g_{0}:=4\sum_{k=1}^{n}\dfrac{\mu}{(|z-z_{k}|^{2}+\mu)^{2}}$ is smooth on $\mathbf{R}^{2}$. With this, we have the following.
Proposition 4.6.
For each $C>0$, a solution of (4.7) is equivalent to a solution $v$ of the following equation
(4.8) $\Delta v-4Ce^{v+u_{0}}-(g_{0}-2C)=0.$
Proof.
Suppose $u$ is a solution of (4.7). Let $v=u-u_{0}$. Then by Lemma 4.5 we have
$\Delta v$ $=\Delta u-\Delta u_{0}$
$=4Ce^{u}-2C+4\pi\sum_{k=1}^{n}\delta(z-z_{k})+g_{0}-4\pi\sum_{k=1}^{n}\delta(z-z_{k})$
$=4Ce^{v+u_{0}}+(g_{0}-2C)$
Thus, $\Delta v-4Ce^{v+u_{0}}-(g_{0}-2C)=0$. The calculations also work backward. ∎
Equation (4.8) is the simplified version of (4.7) that is claimed in Step 1. Once we have the simplified version of our PDE (cf. (4.8)), we move on to Step 2, which shows that (4.8) is the variational equation of a functional. To this end, we define the functional $\mathscr{A}:C^{\infty}_{c}(\mathbf{R}^{2},\mathbf{R})\to\mathbf{R}$ as following (cf. Chap. 3, Sec. 3 in [MR0614447])
(4.9) $\mathscr{A}(v)=\iint_{\mathbf{R}^{2}}\left(\dfrac{1}{2}|\nabla v|^{2}+v(g_{0}-2C)+4Ce^{u_{0}}(e^{v}-1)\right).$
Proposition 4.7.
For each $C>0$, (4.8) is the variational equation of the functional $\mathscr{A}$ defined in (4.9).
Proof.
For each $h\in C^{\infty}_{c}(\mathbf{R}^{2},\mathbf{R})$, the linearization of $\mathscr{A}$ at $v\in C^{\infty}_{c}(\mathbf{R}^{2},\mathbf{R})$ is given by
$d_{v}\mathscr{A}(h)=\displaystyle\lim_{t\to 0}\dfrac{\mathscr{A}(v+th)-\mathscr{A}(v)}{t}.$
We consider the equation $d_{v}\mathscr{A}(h)=0$, for all $h$, which is the equation that finds all critical points of $\mathscr{A}$. Note that
$\mathscr{A}(v+th)=\dfrac{1}{2}\left\lVert\nabla v\right\rVert^{2}$ $+t\langle\nabla v,\nabla h\rangle_{L^{2}}+\displaystyle\iint_{\mathbf{R}^{2}}\dfrac{t^{2}}{2}|\nabla h|^{2}+$
$+\langle v,g_{0}-2C\rangle_{L^{2}}+\iint_{\mathbf{R}^{2}}th(g_{0}-2C)+\iint_{\mathbf{R}^{2}}4Ce^{u_{0}}(e^{v+th}-1).$
As a result, $\dfrac{1}{t}(\mathscr{A}(v+th)-\mathscr{A}(v))$ is equal to
$\langle\nabla v,\nabla h\rangle_{L^{2}}+\displaystyle\iint_{\mathbf{R}^{2}}\dfrac{t}{2}|\nabla h|^{2}+\langle g_{0}-2C,h\rangle_{L^{2}}+\iint_{\mathbf{R}^{2}}4Ce^{u_{0}+v}\dfrac{e^{th}-1}{t}$
$=\langle\nabla v,\nabla h\rangle_{L^{2}}+\displaystyle\iint_{\mathbf{R}^{2}}\dfrac{t}{2}|\nabla h|^{2}+\langle g_{0}-2C,h\rangle_{L^{2}}+\langle 4Ce^{u_{0}+v},h\rangle_{L^{2}}+$
$+\iint_{\mathbf{R}^{2}}4Ce^{u_{0}+v}\sum_{k=2}^{\infty}\dfrac{t^{k-1}h^{k}}{k!}.$
Thus, as $t\to 0$, we have $d_{v}\mathscr{A}(h)=\langle\nabla v,\nabla h\rangle_{L^{2}}+\langle g_{0}-2C+4Ce^{u_{0}+v},h\rangle_{L^{2}}$. Since $\nabla^{*}\nabla=-\Delta$, $d_{v}\mathscr{A}(h)=0$ for all $h$ implies that $-\Delta v+4Ce^{u_{0}+v}+(g_{0}-2C)=0$. ∎
4.2. Some Estimates for $\mathscr{A}$
In this subsection, we deal with Step 3 which was laid out in the previous subsection. We used the term functional rather liberally up to this point. To be more precise, a functional is an $\mathbf{R}$-valued map defined on a certain Banach space such that its absolute value evaluated at each point in the Banach space is finite. Even though, $C^{\infty}_{c}(\mathbf{R}^{2},\mathbf{R})$ is an infinite dimensional vector space, it is not complete. Furthermore, there is no guarantee that $|\mathscr{A}(v)|<\infty$ for all $v\in C^{\infty}_{c}$. To this end, we consider the $L^{2}_{1}$-Sobolev completion of $C^{\infty}_{c}(\mathbf{R}^{2},\mathbf{R})$, where $L^{2}_{1}$ is the Sobolev space of functions that along with their first derivative are in $L^{2}(\mathbf{R}^{2})$. For brevity, when no confusion arises, we simply refer to this completion as $L^{2}_{1}$. Our first task is to show that $\mathscr{A}$ is well-defined on $L^{2}_{1}$.
Proposition 4.8.
For some finite number $c$ and each $v\in L^{2}_{1}$, we have $|\mathscr{A}(v)|\leq e^{c\left\lVert v\right\rVert^{2}_{L^{2}_{1}}}$. In other words, $\mathscr{A}:L^{2}_{1}\to\mathbf{R}$ is well-defined.
Proof.
The proof is similar to the proof of Proposition 3.3 of Ch. 3, Sec. 3 in [MR0614447], but there is a notable difference which we will explain. Note that we can estimate (4.9) as follows
(4.10) $|\mathscr{A}(v)|\leq\dfrac{1}{2}\left\lVert\nabla v\right\rVert^{2}_{L^{2}}+2C\left\lVert v\right\rVert_{L^{1}}+|\langle v,g_{0}+4C(e^{u_{0}}-1)\rangle_{L^{2}}|+4C|\langle e^{u_{0}},e^{v}-v-1\rangle_{L^{2}}|.$
By definition, $\left\lVert\nabla v\right\rVert^{2}_{L^{2}}\leq\left\lVert v\right\rVert^{2}_{L^{2}_{1}}$. For the second term, without loss of generality, we suppose $v$ is a function with compact support. Then by the Hölder inequality, there is a positive constant $c(v)>0$ and some large $c>0$ such that
$\left\lVert v\right\rVert_{L^{1}}\leq c(v)\cdot\left\lVert v\right\rVert_{L^{2}}\leq e^{c\left\lVert v\right\rVert^{2}_{L^{2}}}\leq e^{c\left\lVert v\right\rVert^{2}_{L^{2}_{1}}}.$
Furthermore, by the Cauchy-Schwarz inequality and Proposition 2.3 of Ch. 6 in [MR0614447], we also have
$\langle v,g_{0}+4C(e^{u_{0}}-1)\rangle_{L^{2}}$ $\leq\left\lVert v\right\rVert_{L^{2}}\cdot(\left\lVert g_{0}\right\rVert_{L^{2}}+4C\left\lVert e^{u_{0}}-1\right\rVert_{L^{2}})$
$\leq\left\lVert v\right\rVert_{L^{2}_{1}}\cdot(\left\lVert g_{0}\right\rVert_{L^{2}}+4C\left\lVert e^{u_{0}}-1\right\rVert_{L^{2}}),$
$\langle e^{u_{0}},e^{v}-v-1\rangle_{L^{2}}$ $\leq\left\lVert e^{u_{0}}\right\rVert_{L^{\infty}}\cdot\left\lVert e^{v}-v-1\right\rVert_{L^{1}}\leq\exp(c\left\lVert v\right\rVert^{2}_{L^{2}_{1}}).$
Since $u_{0},g_{0}\in L^{2}$ by Lemma 3.4 of Ch. 3, Sec. 3 in [MR0614447], the proof is complete if we can show that $e^{u_{0}}-1\in L^{2}$. The rest of the proof is to verify this assertion. To this end, we consider the function $f(t)=1-e^{t}+t$ which is non-positive for all $t$. Then $f(u_{0})\leq 0$, which implies that $1-e^{u_{0}}\leq-u_{0}$. Since $u_{0}\leq 0$ by definition, $-u_{0},1-e^{u_{0}}\geq 0$.
Let $R>\max|z_{k}|$. Then for any $|z|\geq 2R$, using the inequality $\ln(1+t)\leq t$ for all real number $t$, we have
$-u_{0}$ $=\sum_{k=1}^{n}\log\left(1+\dfrac{\mu}{|z-z_{k}|^{2}}\right)\leq\sum_{k=1}^{n}\log\left(1+\dfrac{\mu}{(|z|-R)^{2}}\right)$
$\leq\sum_{k=1}^{n}\dfrac{\mu}{(|z|-R)^{2}}=\dfrac{n\mu}{(|z|-R)^{2}}.$
Therefore, for all $|z|\geq 2R$, we have
(4.11) $|e^{u_{0}}-1|=1-e^{u_{0}}\leq\dfrac{n\mu}{(|z|-R)^{2}}.$
Using (4.11), we estimate the $L^{2}$-norm of $12e^{u_{0}}-$ 1 for $|z|\geq 2R$ as follows
[…]
The right-hand-side integral is clearly bounded. To deal with $|z|<2R$, we use the fact that $1-e^{u_{0}}\leq 1$ to establish the $L^{2}$-bound of $e^{u_{0}}-1$ inside a disk. Thus, $e^{u_{0}}-1$ is $L^{2}$ on the entire plane as claimed. The proof of the proposition is now complete. ∎
The next proposition shows that under appropriate conditions, $v=\log|\psi_{2}/\sqrt{C}|^{2}-u_{0}\in L^{2}_{1}$. Thus, it is sufficient to study $\mathscr{A}$ on $L^{2}_{1}$.
Proposition 4.9.
For each $C>0$, let $(A_{0},A_{1},\psi_{1},\psi_{2})$ be a smooth solution of (4.1) such that $\psi_{1}$ and $\psi_{2}$ have the same argument and that for every $\epsilon>0$, there exists a $c=c(\epsilon,C,(A_{0},A_{1},\psi_{1},\psi_{2}))$ such that
$C-|\psi_{2}|^{2}\leq c\cdot e^{-(1-\epsilon)|x|}.$
With $u_{0}$ defined as in Lemma 4.5, then $v=\log|\psi_{2}/\sqrt{C}|^{2}-u_{0}\in L^{2}_{1}$.
Proof.
Same as the proof of Theorem 3.2 of Ch., Sec. 3 in [MR0614447]. ∎
There is a refinement of Proposition 4.7 for $L^{2}_{1}$. In particular, we have the following.
Theorem 4.10.
For any $v$ and $h$ belong to $L^{2}_{1}$, we have
$|\mathscr{A}(v+h)-\mathscr{A}(v)-d_{v}\mathscr{A}(h)|\leq O(1)\left\lVert h\right\rVert^{2}_{L^{2}_{1}}\cdot\exp(c\left\lVert v\right\rVert^{2}_{L^{2}_{1}}+\left\lVert h\right\rVert^{2}_{L^{2}_{1}}),$
$|d_{v}\mathscr{A}(h)|\leq\left\lVert h\right\rVert_{L^{2}_{1}}\exp(c\left\lVert v\right\rVert^{2}_{L^{2}_{1}}).$
Thus, $\mathscr{A}:L^{2}_{1}\to\mathbf{R}$ is differentiable and $d_{v}\mathscr{A}$ defines a bounded linear functional on $L^{2}_{1}$.
Proof.
We use the formula for $d_{v}\mathscr{A}(h)$ derived in the proof of Proposition 4.7 and follow the same proof of Proposition 3.5 of Ch.3, Sec.3 in [MR0614447] to obtain the desired estimates. ∎
To end this subsection, we give an estimate for $d_{v}\mathscr{A}(v)$, which will be relevant for Step 4, where we need to show that $\mathscr{A}:L^{2}_{1}\to\mathbf{R}$ achieves a minimum at a critical point.
Lemma 4.11.
Recall that we have set $\mu>2n/C$. For all $z\in\mathbf{C}$, we have $g_{0}+4Ce^{u_{0}}\leq 2C$. Furthermore, there is a positive constant $\kappa$ such that $2C-g_{0}\geq\kappa$.
Proof.
We prove the first assertion. Note that
$g_{0}+4Ce^{u_{0}}$ $=4\sum_{k=1}^{n}\dfrac{\mu}{(|z-z_{k}|^{2}+\mu)^{2}}+4C\prod_{k=1}^{n}\dfrac{1}{1+\dfrac{\mu}{|z-z_{k}|^{2}}}$
$=4\sum_{k=1}^{n}\dfrac{\mu}{(|z-z_{k}|^{2}+\mu)^{2}}+4C\prod_{k=1}^{n}\dfrac{|z-z_{k}|^{2}}{|z-z_{k}|^{2}+\mu}$
$=\dfrac{4}{\mu}\sum_{k=1}^{n}\dfrac{\mu^{2}}{(|z-z_{k}|^{2}+\mu)^{2}}+4C\prod_{k=1}^{n}\left(1-\dfrac{\mu}{|z-z_{k}|^{2}+\mu}\right).$
Let $\mu_{k}:=\dfrac{\mu}{|z-z_{k}|^{2}+\mu}$. Substituting into the equation, we now have
$\dfrac{4}{\mu}\sum_{k=1}^{n}\dfrac{\mu^{2}}{(|z-z_{k}|^{2}+\mu)^{2}}+4C\prod_{k=1}^{n}\left(1-\dfrac{\mu}{|z-z_{k}|^{2}+\mu}\right)$ $=\dfrac{4n}{\mu}\cdot\dfrac{1}{n}\sum_{k=1}^{n}\mu_{k}^{2}+2C\prod_{k=1}^{n}(1-\mu_{k})$
$\leq 2C\left(\dfrac{1}{n}\sum_{k=1}^{n}\mu_{k}^{2}+\prod_{k=1}^{n}(1-\mu_{k})\right).$
By definition, we always have $0\leq 1-\mu_{k}\leq 1$. So, when applying the inequality $\prod x_{k}\leq(\sum x_{k})/n$ for all $x_{k}\in[0,1]$ to $x_{k}:=\mu_{k}$, we can estimate the right-hand side of the above expression by
$g_{0}+4Ce^{u_{0}}\leq 2C\left(\dfrac{1}{n}\sum\mu_{k}^{2}+\dfrac{1}{n}\sum(1-\mu_{k})\right)\leq 2C\left(1-\dfrac{1}{n}\sum\mu_{k}(1-\mu_{k})\right)\leq 2C.$
The second assertion follows automatically from the fact that $0<g_{0}\leq 4n/\mu<2C$. ∎
Lemma 4.12.
There is a positive finite number $c>0$ and a constant $\kappa>0$ such that for all $v\in L^{2}_{1}$,
$d_{v}\mathscr{A}(v)-\left\lVert\nabla v\right\rVert^{2}_{L^{2}}\geq\kappa\iint_{\mathbf{R}^{2}}\dfrac{v^{2}}{1+|v|}-c.$
Proof.
By the formula derived in the proof of Proposition 4.7, we have
(4.12) $d_{v}\mathscr{A}(v)=\left\lVert\nabla v\right\rVert^{2}_{L^{2}}+\iint_{\mathbf{R}^{2}}v(g_{0}-2C+4Ce^{u_{0}+v}).$
The proof is complete if we can show that the second term on the right-hand side of the above expression is greater or equal to the right-hand side of the claim for some appropriate $\kappa$ and $c$. To do this, we first re-write $v=v^{+}-v^{-}$ into its positive and negative parts. Then we can also re-write the second term on the right-hand-side of (4.12) as
$\underbrace{\iint_{\mathbf{R}^{2}}v^{+}(4Ce^{u_{0}+v^{+}}+g_{0}-2C)}_{\text{$:=I^{+}$}}+\underbrace{\iint_{\mathbf{R}^{2}}(-v^{-})(4Ce^{u_{0}-v^{-}}+g_{0}-2C)}_{\text{$:=I^{-}$}}.$
We estimate each term individually. Observe that
$I^{-}$ $=\iint_{\mathbf{R}^{2}}v^{-}(2C-g_{0}-4Ce^{u_{0}-v^{-}})$
$=\iint_{\mathbf{R}^{2}}v^{-}(2C-g_{0}-4Ce^{u_{0}}+4Ce^{u_{0}}(1-e^{-v^{-}}))$
$\geq\iint_{\mathbf{R}^{2}}v^{-}\left(2C-g_{0}-4Ce^{u_{0}}+4Ce^{u_{0}}\cdot\dfrac{v^{-}}{1+v^{-}}\right)$
$=\iint_{\mathbf{R}^{2}}\dfrac{v^{-}}{1+v^{-}}\left((1+v^{-})(2C-g_{0}-4Ce^{u_{0}})+4Ce^{u_{0}}\cdot v^{-}\right)$
$=\iint_{\mathbf{R}^{2}}\dfrac{v^{-}}{1+v^{-}}\cdot((2C-g_{0}-4Ce^{u_{0}})+(2C-g_{0})v^{-}).$
Here, the third inequality is an application of the estimate $1-e^{-t}\geq t/(1+t)$ for all $t\geq 0$. By Lemma 4.11, we already know that $2C-g_{0}-4Ce^{u_{0}}\geq 0$ and $2C-g_{0}\geq\kappa$. Thus, we have
(4.13) $I^{-}\geq\kappa\iint_{\mathbf{R}^{2}}\dfrac{(v^{-})^{2}}{1+v^{-}}.$
To estimate $I^{+}$, we note that
$4Ce^{u_{0}+v^{+}}+g_{0}-2C$ $=2C(2e^{u_{0}+v^{+}}-1-(u_{0}+v^{+}))+(2Cu_{0}+g_{0}+2Cv^{+})$
$\geq 2Cu_{0}+g_{0}+2Cv^{+}.$
The above inequality is an application of the estimate $2e^{t}-1-t\geq 0$ for all $t\in\mathbf{R}$. As a result,
$I^{+}$ $\geq\iint_{\mathbf{R}^{2}}2C(v^{+})^{2}+v^{+}(2Cu_{0}+g_{0})$
$=\iint_{\mathbf{R}^{2}}2C\left(\dfrac{v^{+}}{2}+\left(u_{0}+\dfrac{g_{0}}{2C}\right)\right)^{2}+\dfrac{3C}{2}(v^{+})^{2}-2C\left(u_{0}+\dfrac{g_{0}}{2C}\right)^{2}$
$\geq\dfrac{3C}{2}\iint_{\mathbf{R}^{2}}(v^{+})^{2}-4C(\left\lVert u_{0}\right\rVert^{2}_{L^{2}}+\left\lVert g_{0}/2C\right\rVert^{2}_{L^{2}})$
(4.14) $\geq\dfrac{3C}{2}\iint_{\mathbf{R}^{2}}\dfrac{(v^{+})^{2}}{1+v^{+}}-4C(\left\lVert u_{0}\right\rVert^{2}_{L^{2}}+\left\lVert g_{0}/2C\right\rVert^{2}_{L^{2}}).$
Since $u_{0},g_{0}\in L^{2}$, combine (4.13) and (4.2) we obtain the desired estimate. ∎
Theorem 4.13.
For all $v\in L^{2}_{1}$, there a constant $\epsilon>0$ and a finite positive number $c$ such that
$d_{v}\mathscr{A}(v)\geq\epsilon\left\lVert v\right\rVert_{L^{2}_{1}}-c.$
Proof.
By Lemma 4.12, there are positive finite constants $\kappa,c$ such that
$d_{v}\mathscr{A}(v)$ $\geq\left\lVert\nabla v\right\rVert^{2}_{L^{2}}+\kappa\iint_{\mathbf{R}^{2}}\dfrac{v^{2}}{1+|v|}-c$
$=\left\lVert v\right\rVert^{2}_{L^{2}_{1}}-\left\lVert v\right\rVert^{2}_{L^{2}}+\kappa\iint_{\mathbf{R}^{2}}\dfrac{v^{2}}{1+|v|}-c$
(4.15) $\geq\left\lVert v\right\rVert^{2}_{L^{2}_{1}}(1-\lambda(v))+\dfrac{\kappa}{2}\cdot\dfrac{\lambda^{2}(v)\left\lVert v\right\rVert^{2}_{L^{2}_{1}}}{1+\left\lVert v\right\rVert_{L^{2}_{1}}}-c.$
Here, $\lambda(v)$ is a number such that $\left\lVert v\right\rVert^{2}_{L^{2}}=\lambda(v)\left\lVert v\right\rVert^{2}_{L^{2}_{1}}$. Note that $\lambda(v)\in[0,1]$. The estimation of the second term on the right-hand-side of (4.2) is given by the Cauchy-Schwarz inequality and the Sobolev embedding theorem as follows
$\lambda^{2}(v)\left\lVert v\right\rVert^{4}_{L^{2}_{1}}$ $=\left(\iint_{\mathbf{R}^{2}}v^{2}\right)^{2}$
$=\left(\iint_{\mathbf{R}^{2}}\dfrac{v}{\sqrt{1+|v|}}\cdot v\sqrt{1+|v|}\right)^{2}$
$\leq\iint_{\mathbf{R}^{2}}\dfrac{v^{2}}{1+|v|}\cdot\iint_{\mathbf{R}^{2}}(v^{2}+|v|^{3})=\iint_{\mathbf{R}^{2}}\dfrac{v^{2}}{1+|v|}\cdot(\left\lVert v\right\rVert^{2}_{L^{2}}+\left\lVert v\right\rVert^{3}_{L^{3}})$
$\leq\iint_{\mathbf{R}^{2}}\dfrac{v^{2}}{1+|v|}\cdot\left(\left\lVert v\right\rVert^{2}_{L^{2}_{1}}+\left\{\pi\left(\dfrac{3-2}{2}\right)\right\}^{\dfrac{3-2}{2}}\left\lVert v\right\rVert^{3}_{L^{2}_{1}}\right)$
$\leq\iint_{\mathbf{R}^{2}}\dfrac{v^{2}}{1+|v|}\cdot 2\left\lVert v\right\rVert^{2}_{L^{2}_{1}}(1+\left\lVert v\right\rVert_{L^{2}_{1}}).$
Now we view the right-hand-side of (4.2) as a quadratic function of $\lambda$-variable, denoted by $h(\lambda)$, where $\lambda\in[0,1]$. The critical point of $h(\lambda)$ is given by $\lambda_{0}=(1+\left\lVert v\right\rVert_{L^{2}_{1}})/\kappa$. There are two cases to consider:
Case 1: $\lambda_{0}\leq 1$. Then $h(\lambda)$ achieves minimum at $\lambda_{0}$. Denote $t=\left\lVert v\right\rVert_{L^{2}_{1}}$ and note that $10\leq t\leq\kappa-$ 1. The right-hand-side of (4.2) is always at least
$g(t):=h(\lambda_{0})=-\dfrac{1}{\kappa}t^{3}+\left(\dfrac{\kappa-1}{\kappa}+\dfrac{1}{2\kappa}\right)t^{2}+\dfrac{1}{2\kappa}t-c.$
As $g(t)$ is a cubic function in $t$-variable, on $[0,\kappa-1]$ one can always find a linear function $\epsilon t-c^{\prime}$, where $\epsilon,c^{\prime}>0$ such that $h(\lambda(v))\geq g(t)\geq\epsilon t-c^{\prime}=\epsilon\left\lVert v\right\rVert_{L^{2}_{1}}-c^{\prime}$.
Case 2: $\lambda_{0}>1$. Then $h(\lambda)$ achieves minimum at $\lambda=1$. In other words, $h(\lambda(v))\geq h(1)$, where $h(1)$ is computed by
$\dfrac{\kappa}{2}\cdot\dfrac{\left\lVert v\right\rVert^{2}_{L^{2}_{1}}}{1+\left\lVert v\right\rVert_{L^{2}_{1}}}-c:=\dfrac{\kappa}{2}\cdot\dfrac{t^{2}}{1+t}-c.$
Since $\lambda_{0}>1$, $t>\kappa-1$. We pick $\tau$ to be a positive constant such that $\tau<(\kappa-1)/\kappa<1$. Then we always have $t^{2}\geq\tau t(1+t)$, where $t>\kappa-1$. As a result, we also have $h(\lambda(v))\geq\tau\left\lVert v\right\rVert_{L^{2}_{1}}-c$. ∎
4.3. Infinimum of $\mathscr{A}$
We proceed to Step 4 and Step 5 which were laid out in Subsection 4.1. Recall that we work to show the variational equation of $\mathscr{A}$ (cf. (4.8), Proposition 4.7) has a solution. If (4.8) has a solution, then by Proposition 4.6, (4.7) also has a solution. This in turn implies that the Seiberg-Witten vortex equation without Higgs fields (4.1) has a solution that has exponential decay property (property (E)) and the $U(1)$-connection part of the solution is not necessarily flat (cf. Theorem 4.1). Note that if a functional defined on a Banach space is differentiable around its local extrema, then such an extrema also satisfies the variational equation of the functional (cf. Prop 7.5 of Ch. 6, Sec. 7 in [MR0614447]). So, we prove that $\mathscr{A}$ has extrema.
Lemma 4.14.
The functional (4.9) $\mathscr{A}:L^{2}_{1}\to\mathbf{R}$ is strictly convex, i.e, for any $v,w\in L^{2}_{1}$ and $t\in(0,1)$, we have $\mathscr{A}((1-t)v+tw)<(1-t)\mathscr{A}(v)+t\mathscr{A}(w)$.
Proof.
It is sufficient to show that the functional $v\mapsto e^{v}-v-1$ is strictly convex. But this is obvious from the fact that the function $e^{t}-t-1$ is strictly convex on the entire real line. ∎
Proposition 4.15.
The functional (4.9) $\mathscr{A}:L^{2}_{1}\to\mathbf{R}$ is weakly lower semi-continuous on $L^{2}_{1}$. By weakly lower semi-continuous, we mean that if $v_{k}\rightharpoonup v$ weakly in $L^{2}_{1}$, then $\mathscr{A}(v_{k})\geq\mathscr{A}(v)$.
Proof.
Let $v$ be the weak limit of a sequence $\{v_{k}\}$ in $L^{2}_{1}$. By Lemma 4.14, $\mathscr{A}$ is strictly convex and we have
$\mathscr{A}(v+t(v_{k}-v))<(1-t)\mathscr{A}(v)+t\mathscr{A}(v_{k}).$
After re-arranging the above inequality and let $t\to 0$ (which is possible because of Theorem 4.10), we obtain
$d_{v}\mathscr{A}(v_{k}-v)=\lim_{t\to 0}\dfrac{\mathscr{A}(v+t(v_{k}-v))-\mathscr{A}(v)}{t}\leq\mathscr{A}(v_{k})-\mathscr{A}(v).$
A bounded linear functional is also weakly continuous. By Theorem 4.10, $d_{v}\mathscr{A}$ is bounded. Therefore, $\mathscr{A}(v)\leq\displaystyle\lim_{k\to\infty}\mathscr{A}(v_{k})$ as claimed. ∎
Theorem 4.13 in the previous subsection tells us that for all $v\in L^{2}_{1}$, there are positive finite constants $\epsilon,c$ such that $d_{v}\mathscr{A}(v)\geq\epsilon\left\lVert v\right\rVert_{L^{2}_{1}}-c$. Thus for $R>c/\epsilon$ and any $\left\lVert v\right\rVert_{L^{2}_{1}}=R$, $\inf_{\left\lVert v\right\rVert_{L^{2}_{1}}=R}d_{v}\mathscr{A}(v)\geq\epsilon\cdot R-c>0$. Furthermore, by Proposition 4.15 and Theorem 4.10, $\mathscr{A}$ is weakly lower semi-continuous and strongly differentiable. Proposition 8.6 of Ch. 6, Sec. 8 in [MR0614447] tells us that $\mathscr{A}$ has a local minimum in the ball $\left\lVert v\right\rVert_{L^{2}_{1}}<R$. Thus, we have completed Step 4.
Theorem 4.16.
The functional (4.9) $\mathscr{A}:L^{2}_{1}\to\mathbf{R}$ has a local minimum.
To complete the proof of Theorem 4.4, we move on to Step 5, which shows that the solution $u$ of (4.7) is real analytic. It is sufficient to show the following
Proposition 4.17.
The solution $v\in L^{2}_{1}$ of (4.8) is real analytic on the entire plane.
Proof.
The proof is a standard bootstrapping argument and the fact that solutions to elliptic equations are real analytic. ∎
4.4. Zeroes of Exponential Decay Solutions
In Subsection 4.1, we explicitly use the fact that the functions $\psi_{1}$ and $\psi_{2}$ have only a finite zero set. We justify this assumption with the following three lemmas that appeal to properties of the functions.
Lemma 4.18.
Let $f$ be a complex-valued function defined on $\mathbf{C}$ such that there exist positive constants $M,N$ and
$0\leq M-|f(z)|^{2}\leq Ne^{-|z|},$
for every $z$. There exists a radius $r>0$ so that every zero of the function $f$ is contained in the closed disk of radius $r$.
Proof.
Suppose for every $R>0$ there exists a $z_{R}$ such that $|z_{R}|>R$ and $f(z_{R})=0$. Observe that
$0<M\leq Ne^{-|z_{R}|}\leq Ne^{-R},$
for every $z_{R}$ and $R$. Since such a $z_{R}$ exists for every $R>0$, it follows that
$0=\lim_{R\to\infty}Ne^{-R}\geq\lim_{R\to\infty}Ne^{-|z_{R}|}\geq M>0,$
which is a contradiction. ∎
Lemma 4.19.
Let $A,B\in L^{q}(D)$, $q>2$. Every solution $w$ of the Vekua equation
$\frac{\partial w}{\partial\overline{z}}=Aw+B\overline{w}$
has the form $w=he^{\gamma}$, where $e^{\gamma}$ is Hölder continuous and $h$ is a holomorphic function. A complex number $z$ is a zero of $w$ if and only if $z$ is a zero of $h$.
Proof.
By Theorem 2.4, every solution of the Vekua equation has the form $w=he^{\gamma}$, where $h$ is holomorphic and $e^{\gamma}$ is Hölder continuous. Since $|e^{\gamma(z)}|>0$, for all $z$, it follows that if $w(z)=0$, then $h(z)=0$. In the other direction, if $h(z)=0$, then $w(z)=h(z)e^{\gamma(z)}=0$. ∎
Lemma 4.20.
Every function $f$ that is holomorphic on a closed disk and not identically zero has a finite zero set.
Proof.
Suppose that the zero set of $f$ is infinite. Since the closed disk is compact, it follows that there is a subsequence that converges to a point of the closed disk by the Bolzano-Weierstrass Property [Folland]. This implies that $f$ is the identically zero function [GK], which is a contradiction. ∎
We combine the preceding three lemmas to justify the following proposition.
Proposition 4.21.
Let $A,B\in L^{q}(D)$, $q>2$. Every function $w$ that solves
$\frac{\partial w}{\partial\overline{z}}=Aw+B\overline{w}$
on $\overline{D}$ such that there exist positive constants $M,N$ that satisfy
$0\leq M-|w(z)|^{2}\leq Ne^{-|z|}$
has a finite zero set.
The last proposition allows us to characterize the zero sets of the $\psi_{1},\psi_{2}$ that are components of solutions to (4.1) by associating them with the systems considered in Section 2.
Theorem 4.22.
Let $A\in L^{q}(D)$, $q>2$. Every solution pair $(w_{1},w_{2})$ of the system
$\begin{cases}\dfrac{\partial w_{1}}{\partial z}+iAw_{1}+Bw_{2}=0\\ &\\ \dfrac{\partial w_{2}}{\partial\overline{z}}+i\overline{A}w_{2}+\overline{B}w_{1}=0,\end{cases}$
such that $B\dfrac{w_{2}}{w_{1}},\overline{B}\dfrac{w_{1}}{w_{2}}\in L^{q}(D)$, $q>2$, and there exist positive constants $M_{1},N_{1},M_{2},N_{2}$ such that
$0\leq M_{1}-|w_{1}(z)|^{2}\leq N_{1}e^{-|z|}$
and
$0\leq M_{2}-|w_{2}(z)|^{2}\leq N_{2}e^{-|z|},$
for every $z\in\mathbf{C}$, has a finite zero set.
Proof.
Recall from Theorem 2.6 that solutions of this type of system have the form $(w_{1},w_{2})=(\overline{e^{\gamma_{1}}}\overline{h_{1}},e^{\gamma_{2}}h_{2})$, where the $e^{\gamma_{j}}$ are Hölder continuous functions and the $h_{j}$ are holomorphic, $j=1,2$. The result follows immediately for $w_{2}$ by Proposition 4.21. The result follows for $w_{1}$ by applying Proposition 4.21 to $\overline{w_{1}}$ and recognizing that it has the same zero set as $w_{1}$. ∎
An immediate corollary of the above theorem that is relevant in our situation is the following.
Corollary 4.23.
If $(A_{0},A_{1},\psi_{1},\psi_{2})$ is a solution of the Seiberg-Witten vortex equation (4.1) that satisfies the property (E), i.e., for some positive number $C>0$, $\psi_{1},\psi_{2}$ have the same argument and that for every $\epsilon>0$, there exists a $c>0$ such that $C-|\psi_{2}|^{2}\leq c\cdot\exp(-(1-\epsilon)|z|)$, then $\psi_{1},\psi_{2}$ have a finite set of zero.
This concludes the proof of Theorem 4.1 (cf. Theorem 1.3).
5. Discussion
The proof of Theorem 4.1 (cf. Theorem 1.3) relies on the algebraic procedure of turning a system of non-linear PDEs (cf. (4.1)) into a single Kazdan-Warner-type equation (cf. (4.7), (4.8)). As a result, the solutions of (4.7) or (4.8) that exponentially go to zero characterize solutions with exponential decay of the Seiberg-Witten vortex equations. For the Seiberg-Witten vortex equations with Higgs fields (cf. (3.12))
$\begin{cases}\dfrac{\partial\phi}{\partial\overline{z}}=-\dfrac{1}{2}\psi_{1}\overline{\psi_{2}},\\ i\left(\dfrac{\partial A_{1}}{\partial x_{0}}-\dfrac{\partial A_{0}}{\partial x_{1}}\right)=\dfrac{i}{2}(|\psi_{1}|^{2}-|\psi_{2}|^{2})\\ 2\dfrac{\partial\psi_{2}}{\partial\overline{z}}+i(A_{0}+iA_{1})\psi_{2}-\overline{\phi}\psi_{1}=0\\ 2\dfrac{\partial\psi_{1}}{\partial z}+i(A_{0}-iA_{1})\psi_{1}-\phi\psi_{2}=0,\end{cases}$
the elimination process of unknown variables is much more cumbersome and does not yield a single PDE that characterizes the solutions of (3.12) in the same streamlined manner as in the case of the Seiberg-Witten vortex equations without Higgs fields. Note that in [MR1952133], it is shown that the moduli space of (3.12) is non-empty.
Proposition 5.1 (cf. Proposition 2.2 in [MR1952133]).
Let $A=iA_{0}dx_{0}+iA_{1}dx_{1}$ (re-written in complex coordinate as $A=A^{1,0}dz+A^{0,1}d\overline{z}$). The moduli space of (3.12) contains non-trivial solutions. Specifically, for any $c_{1}\in\mathbf{C}$ and $c_{2}\in\mathbf{R}$ such that $|c_{1}|=\sqrt{2}c_{2}$, then
(5.1) $(A,\psi_{1},\psi_{2},\phi)=\left(\dfrac{-ic_{2}}{2}dz,c_{1},c_{1}e^{ic_{2}(z+\overline{z})},-ic_{2}e^{-ic_{2}(z+\overline{z})}\right)$
is always a solution of (3.12).
One should compare the above with our Proposition 3.9. The solutions of the type (5.1) have no zeroes (unless $c_{1},c_{2}$ are zero of course) and the connection $A$ is always flat. It is not difficult to see that to obtain zeroes, one only needs to re-scale by appropriate polynomial function in either $z$ or $\overline{z}$ variable. As a result, all of the solutions of (3.12) of this type have polynomial growth. In light of our Theorem 4.1, even though the method of analysis in this paper cannot be applied directly to the situation of (3.12), we make the following conjecture.
Conjecture 5.2.
There exists smooth solution $(A,\psi_{1},\psi_{2},\phi)$ of (3.12) such that $\psi_{1},\psi_{2}$ satisfy the following property: For some positive real constants $C_{1},C_{2}$, we have
$0\leq C_{1}-|\psi_{1}|^{2}\leq O(\exp(-|z|)),\quad\quad 0\leq C_{2}-|\psi_{2}|^{2}\leq O(\exp(-|z|)).$
In other words, we expect that the solutions of (3.12) can exhibit both polynomial growth and exponential decay behavior.
Recall that the Seiberg-Witten vortex equations (with or without Higgs fields) are derived from the Seiberg-Witten equations in dimension four. There are other generalizations of the Seiberg-Witten equations that have been considered recently (see e.g., [Taubes:2016voz, MR2980921, MR1392667, haydys2015compactness, nguyen2023pin, sadegh2023threedimensionalseibergwittenequations32spinors, sadegh2024fueteroperator32spinors]). The dimensional reduction employed in this paper is not the naïve one where we simply forget the other dimensions. As a result, the analysis of the moduli space in the 2-dimensional theory still lends some insights about the moduli space in higher dimensions. Therefore, it is interesting to investigate the same analysis question that is considered in this paper for the dimensional reduction of the many generalizations of the Seiberg-Witten equations. We will address this direction of research in the future.
Additionally, the systems of Vekua-type equations considered in Section 2 and used in Section 4 to justify properties of the zero sets of solutions to vortex equations are not only a new structure but arise naturally from the consideration of gauge equations. This indicates that they merit further study of their own.
[…]